### v1

# Quantum Wave Function Collapse for Procedural Content Generation

Abstract
Quantum computers exhibit an inherent randomness, so it seems natural to consider them for procedural content generation. In this work, a quantum version of the famous (classical) wave function collapse algorithm is proposed. This quantum wave function collapse algorithm is based on the idea that a quantum circuit can be prepared in such a way that it acts as a special-purpose random generator for content of a desired form. The proposed method is presented theoretically and investigated experimentally on simulators and actual IBM Quantum devices.
Wave function collapse (WFC) is a powerful tool for procedural content generation (PCG) that is for example used in video games, as it can save significant development time by automating the creation of diverse and complicated game elements. However, it is not limited to video games, but also has applications in various other fields, including art and design, where the need for algorithmically generated content is widespread. Content in this context can therefore have a broad range of meanings: Images, 3D models, game levels, text, sound or a combination of these, to name just a few examples.
Originally proposed in Ref. [1], the WFC algorithm is a a non-backtracking, greedy constraint solving method [2, 3] that is able to generate complex patterns based on a set of input samples. It is known for its ability to create diverse and complex outputs that resemble the input samples while exhibiting novel combinations and variations. WFC employs two implementation strategies: the simple tiled model and the overlapping model, which share an identical algorithm core [4]. In the simple tiled model, tilesets are manually prescribed with predefined adjacency constraints, whereas the overlapping model automatically generates this information from a sample input.
In both strategies, the output is compartmentalized into segments and the possibilities for each segment are iteratively constrained until a unique solution is determined. The term “wave function collapse” is borrowed from quantum physics because of the conceptual similarity. The “wave function” refers to the set of potential states of the segments, whereas the ”collapse“ occurs during the iterative process of narrowing down the possibilities.
Despite its name, WFC is a purely classical algorithm. But even if WFC has nothing to do with quantum physics beyond the terminology, the question naturally arises as to whether quantum computers can be used to execute a genuine quantum version of the algorithm. The investigation of this question and a first draft for a corresponding algorithm are the subject of this manuscript. For this purpose, the terms classical wave function collapse (CWFC), quantum wave function collapse (QWFC), and hybrid quantum-classical wave function collapse (HWFC) are used In the following to clearly distinguish between classical, quantum and hybrid quantum-classical WFC.
Leveraging quantum computers for PCG has been considered before, mainly focusing on the quantum generalization of a blurring process [5, 6] and map generation using a quantum-enhanced decision making process [7]. Related projects have been carried out in connection with the development of quantum-inspired games [8].
The contribution of this paper is threefold:
  • A probabilistic formulation of the simple tiled model of CWFC is presented as a foundation for a quantum version.
  • A QWFC method and a HWFC method are proposed.
  • The proposed methods are tested on simulators and actual IBM Quantum devices.
The remaining manuscript is structured as follows. In the next section, a formal description of CWFC is provided in an appropriate probabilistic form, which enables the development of a QWFC method and a HWFC method in the subsequent section. In the section after, the proposed methods are demonstrated in practice. Finally, the paper ends with a conclusion and an outlook.
1 CLASSICAL METHODS
In this section, the formal description for a probabilistic iterative PCG (PIPCG) is provided and, subsequently, it is shown how it can be reduced to the common CWFC approach.
1.1 PIPCG
To begin with, it is first necessary to formally define the generated content at a sufficiently abstract level. For this purpose, it is presumed that any content instance $C$ can be described as an ordered sequence of $N$ segments
$C:=\{(1,v_{1}),\dots,(N,v_{N})\}\in\mathcal{C},$ (1)
where each segment $(i,v_{i})$ is defined by a unique identifier $i\in[1,N]$ and a value $v_{i}\in A_{i}$ from an alphabet $A_{i}\subset\mathbb{N}$. Here and in the following, the notation $[i,j]:=\{i,\dots,j\}\,\forall\,i,j\in\mathbb{N}$ is used. The set of all possible content instances $\mathcal{C}$ is determined by the content alphabets $\mathcal{A}:=\{A_{1},\dots,A_{N}\}$ for a given size $N$ with $|\mathcal{C}|=\prod_{i=1}^{N}A_{i}$. For example, if the content represents an image, its individual pixels can constitute the segments. Each pixel can attain one color from the available color space of the image (as its alphabet).
1.1.1 Iterative process.
The procedural generation of content instance $C\in\mathcal{C}$ is realized through an iterative process, where each iteration $k\in[1,N]$ consists of three steps:
  1. 1.
    Randomly choose a new segment identifier $s(k)\in[1,N]$ from a set of available identifiers.
  2. 2.
    Randomly choose a corresponding value $v_{s(k)}\in A_{s(k)}$ from a set of available values.
  3. 3.
    Add the chosen segment $(s(k),v_{s(k)})$ to the content instance.
By repeating these steps, the content instance is assembled in the sense of $C(0)\overset{k=1}{\rightarrow}C(1)\overset{k=2}{\rightarrow}\cdots\overset{k=N}{\rightarrow}C(N)$, where
$C(k):=\{(s(1),v_{s(1)}),\dots,(s(k),v_{s(k)})\}\in\mathcal{C}(k),$ (2)
represents the generated partial content instance at the end of iteration $k$ with $|C(k)|=k$, $C(k)=C(k-1)\cup\{(s(k),v_{s(k)})\}$, $s(i)\neq s(j)\,\forall\,i,j\in[1,k]\land i\neq j$. Here, $\mathcal{C}(k)$ denotes all possible partial content instances at the end of iteration $k$. Thus, $|\mathcal{C}(k)|=\sum_{I\in J(k)}\prod_{i\in I}|A_{i}|$, where $J(k):=\{I\,|\,I\in[1,N]\land|I|=k\}$ with $|J(k)|={N\choose k}$. It is convenient to define $C(0):=\{\}$ and $\mathcal{C}(0):=\{C(0)\}$. In addition,
$S^{+}(C(l)):=\{s(1),\dots,s(l)\}$ (3)
denotes the set of segment indices (dropping the values) that are contained in the partial content instance $C(l)$ with $|S^{+}(C(l))|=l$ and
$S^{-}(C(l)):=[1,N]\setminus S^{+}(C(l))$ (4)
denotes the remaining segment indices that are not contained in $C(l)$ with $|S^{-}(C(l))|=N-l$ for all $C(l)\in\mathcal{C}(l)$ and for all $l\in[0,N]$.
After $N$ iterations, the newly generated content instance is complete, i.e., $C(N)\equiv C$ and $\mathcal{C}(N)\equiv\mathcal{C}$. In the following, the three steps within each iteration $k$ are explained in more detail.
1.1.2 First step.
For the first step, an identifier is selected by drawing a sample $s(k)$ from the random variable
$\boldsymbol{s}(k)\sim p^{\mathrm{s}}_{k}(i=s(k)|C^{\prime}=C(k-1)).$ (5)
The probability distribution $p_{k}^{\mathrm{s}}(i|C^{\prime})$ with support $i\in[1,N]$ and $C^{\prime}\in\mathcal{C}(k-1)$ is a user-defined parameter of the procedure that determines how new segment indices $l$ are selected based on the partial content instance $C^{\prime}=C(k-1)$ that has already been generated up to the previous iteration $k-1$. The dependence on all previously chosen segments makes the process generally non-Markovian. It is required that two conditions hold true for the probability distribution $p_{k}^{\mathrm{s}}(i|C^{\prime})$:
  1. S1
    No identifier can be chosen twice, i.e., $p_{l}^{\mathrm{s}}(i=s|C^{\prime}=C(l-1))=0\,\forall\,s\in S^{+}(C(l-1))\,\forall\,C(l-1)\in\mathcal{C}(l-1)\,\forall\,l\in[1,N]$.
  2. S2
    At least one valid identifier has to be available, i.e., $\forall\,l\in[1,N]\,\forall\,C(l-1)\in\mathcal{C}(l-1)\,\exists\,s\in S^{-}(C(l-1)):p_{l}^{\mathrm{s}}(i=s|C^{\prime}=C(l-1))>0$.
The segment identifier selection mechanism is one of the main ingredients of CWFC. Here, it is presented in very generic way. The common entropic approach will be explained further below along with possible variants that are more suitable for a quantum approach.
1.1.3 Second step.
After the identifier $s(k)$ has been chosen, the second step is to select a corresponding value $v_{s(k)}$ for the segment. For this purpose, a sample $v_{s(k)}$ is drawn from the random variable
$\boldsymbol{v}_{s(k)}\sim p^{\mathrm{v}}_{k}(\nu=v_{s(k)}|i=s(k),C^{\prime}=C(k-1)).$ (6)
The probability distribution $p^{\mathrm{v}}_{k}(\nu|i,C^{\prime})$ with support $\nu\in\cup_{A\in\mathcal{A}}A$, $i\in[1,N]$ and $C^{\prime}\in\mathcal{C}(k-1)$ is a user-defined parameter of the procedure that determines how new values are selected based on the selected identifier $i=s(k)$ of the current iteration and the partial content instance $C^{\prime}=C(k-1)$ that has already been generated up to the previous iteration $k-1$. It is required that two conditions hold true for the probability distribution $p^{\mathrm{v}}_{k}(\nu|i,C^{\prime})$:
  1. V1
    Only values from the corresponding alphabet are available, i.e., $p^{\mathrm{v}}_{l}(\nu=v|i=s,C^{\prime}=C(l))=0\,\forall\,v\not\in A_{s}\,\forall\,s\in[1,N]\,\forall\,C(l)\in\mathcal{C}(l-1)\,\forall\,l\in[1,N]$.
  2. V2
    At least one value from the corresponding alphabet has to be available, i.e., $\forall\,l\in[1,N]\,\forall\,C(l-1)\in\mathcal{C}(l-1)\forall s\in S^{-}(C(l-1))\exists v\in A_{s}:p_{l}^{\mathrm{v}}(\nu=v|i=s,C^{\prime}=C(l))>0$.
As explained further below, it is also possible to specify $p^{\mathrm{v}}_{k}(\nu|i,C^{\prime})$ implicitly based on a set of rules, which is the default approach for CWFC and much more convenient for practical purposes than an explicit definition. To simplify the notation a joint distribution for the segment identifier and value selection $p_{k}(i,\nu|C^{\prime}):=p_{k}^{\mathrm{s}}(i|C^{\prime})p_{k}^{\mathrm{v}}(\nu|i,C^{\prime})$ can be defined.
1.1.4 Third step.
In the third step, the selected tuple $(s(k),v_{s(k)})$ that constitutes the newly generated segment is added to the already generated segments of the content instance, i.e., $C(k)=C(k-1)\cup\{(s(k),v_{s(k)})\}$. If $k<N$, the process is repeated and another segment is generated for the next iteration $k+1$. If $k=N$, the PIPCG is finished.
1.1.5 Instance distribution.
An essential concept of PCG is, that only a limited subset of all possible content instances $C\in\mathcal{C}$ is generated according to the designer’s aesthetic preferences, a given rule set, or some other underlying logic. Otherwise, the instances could be drawn directly from $\mathcal{C}$ with much less effort. For the presented formalism, the probability to generate a partial content instance $C(k)$ in the form of eq. 2 is given by
$p(C(k)):=\sum_{\sigma\in S_{k}(C(k))}\prod_{l=1}^{k}q_{l}^{k}(\sigma,C(k))$ (7)
with the abbreviation
$q_{l}^{k}(\sigma,C(k)):=p_{l}($ $i=\sigma_{l},\nu=v_{\sigma_{l}}|$
$C^{\prime}\!=\!\{(\sigma_{1},v_{\sigma_{1}}),\!\dots,\!(\sigma_{l-1},v_{\sigma_{l-1}})\}).$ (8)
Here, $S_{k}(C(k))$ denotes the set of all permutations of $S^{+}(C(k))$, eq. 3, with $|S_{k}(C(k))|=k!$ and $\sigma_{l}$ represents the $l$th value of a permutation $\sigma\in S_{k}(C(k))$ for $l\in[1,k]$. The notation in section 1.1.5 is to be understood in such a way that $C^{\prime}=\{\}$ for $l=1$. Moreover, $v_{\sigma_{m}}$ in section 1.1.5 represents the value of the segment with identifier $\sigma_{m}$ from the content instance $C(k)$ for $m\in[1,l]$. By definition, the partial content instance $C(k)$ can only be generated if $p(C(k))>0$. In other words, the support
$\underline{\mathcal{C}}(k):=\{C(k)\,|\,C(k)\in\mathcal{C}(k)\,\land\,p(C(k))>0\}$ (9)
of $p(C(k))$ contains all partial content instances that can be generated for the given $k$.
For $k=N$, consider a complete content instance $C$ in the form of eq. 1 with values $v_{1},\dots,v_{N}$. In this case, eq. 7 reads
$p(C)=p(v_{1},\dots,v_{N}):=\sum_{\sigma\in S_{N}}\prod_{l=1}^{N}q_{l}^{N}(\sigma,C),$ (10)
where $S_{N}:=S_{N}(C)$ is the set of all permutations of $[1,N]$. In fact, performing the presented PIPCG procedure corresponds to drawing a sample from this distribution such that the content instance can be considered as a random variable $\boldsymbol{C}\sim p(C)$. Consequently, $\underline{\mathcal{C}}(N)=\underline{\mathcal{C}}:=\{C\,|\,C\in\mathcal{C}\,\land\,p(C)>0\}$ denotes the set of all content instances that can be generated.
1.2 Classical Wave Function Collapse (CWFC)
CWFC is a special case of the previously described PIPCG. In this manuscript, only a simplified form of the simple tiled model of CWFC is considered, as it is sufficient to capture the key components of the approach. A detailed description of the original CWFC implementation can be found in, e.g., Refs. [2, 3].
For the common CWFC approach, a single alphabet is used for all segments, i.e., $A_{i}=A\,\forall\,i\in[1,N]$. Given $W$ symbols, one can presume without loss of generality that
$A:=[1,W].$ (11)
Furthermore, the segment identifier selection is performed with an entropic approach, whereas the value selection is pattern-based. These two concepts will be explained in more detail in the following.
1.2.1 Value selection.
For the pattern-based value selection, the segments are organized in such a way that they have a predefined adjacency relationship to each other that does not change. For example, consider that the pixels of an image constitute the segments. Choosing adjacency between the nearest neighbors leads to an adjacency configuration with four directions (right, up, left, down).
Formally, each direction within a adjacency configuration is specified by an index $d\in[1,D]$, where $D$ denotes the total number of directions. The direction-based adjacency relationship between different segments is then defined by the coefficients $\alpha^{d}_{i,j}\in\mathbb{B}:=\{0,1\}$ for $i,j\in[1,N]$ and $d\in[1,D]$. If $\alpha^{d}_{i,j}=1$, segment $i$ is connected to segment $j$ in direction $d$; otherwise, it is not connected. For each direction $d$, these coefficients form an adjacency matrix $\alpha^{d}\in\mathcal{B}:=\mathbb{B}^{N\times N}$ of a directed graph with $N$ vertices that represent the segments. Hence, the segment adjacency is in fact an abstract concept that is not necessarily related to the visual form of the content. An example is shown in fig. 1.
Figure 1: Exemplary nearest neighbor adjacency configuration for CWFC on a two-dimensional grid consisting of $N=9$ segments. The adjacency relationship for each of the four directions right ($d=1$), up ($d=2$), left ($d=3$), and down ($d=4$) can be represented as a directed graph with adjacency matrix $\alpha^{d}$.
Based on the chosen adjacency configuration, a pattern-based value selection can be defined. For this purpose, a more general rule-based selection is introduced first and then the pattern-based selection is presented as a special case.
For a rule-based value selection, consider a set of $m$ rules
$R:=\{r_{1},\dots,r_{m}\}\in\mathcal{R},$ (12)
from the set of of all possible rulesets $\mathcal{R}$, where each rule
$r_{i}:=r_{i}(v,w):=(v,w)$ (13)
for $i\in[1,m]$ consists of a value $v\in A$ and a weight function $w:=w(i,\alpha^{d},C^{\prime})\in\mathbb{R}_{\geq 0}$ with
$w:[1,N]\times\mathcal{B}\times\overline{\mathcal{C}}\rightarrow\mathbb{R}_{\geq 0},$ (14)
where
$\overline{\mathcal{C}}:=\cup_{k=0}^{N-1}\mathcal{C}(k)$ (15)
denotes the unified set of all partial contents. Given a ruleset $R\in\mathcal{R}$, the rule-based version of the probability distribution from eq. 6 is defined as
$p_{k}^{\mathrm{v}}(\nu|i,C^{\prime})=p_{k}^{\mathrm{v,R}}(\nu|i,C^{\prime}):=\frac{F_{k}(\nu,i,C^{\prime})}{\overline{F}_{k}(i,C^{\prime})}$ (16)
with the abbreviations
$F_{k}(\nu,i,C^{\prime}):=\sum_{\begin{subarray}{c}(v,w)\in R\\ v=\nu\end{subarray}}w(i,\alpha^{d},C^{\prime})$ (17)
and
$\overline{F}_{k}(i,C^{\prime}):=\sum_{\nu\in A}F_{k}(\nu,i,C^{\prime}),$ (18)
respectively, for $\nu\in A$, $i\in[1,N]$, $C^{\prime}\in\mathcal{C}(k)$ for $k\in[1,N]$, and $\alpha^{d}\in\mathcal{B}$ for $d\in[1,D]$. Summarized, the set of rules $R$ determines how values are selected based on the non-vanishing results of the respective weight functions. Note that condition V1 is always satisfied, whereas in case of $\overline{F}_{k}(i,C^{\prime})\overset{!}{=}0$ for any $C^{\prime}\in\mathcal{C}(k)$ with $p_{k}^{\mathrm{s,H}}(i|C^{\prime})p(C^{\prime})>0$, condition V2 is violated.
Pattern-based value selection is a special case of rule-based value selection, where the weight function
$w(i,\alpha^{d},C^{\prime})\!:=\!w^{\mathrm{P}}(i,\alpha^{d},C^{\prime},P,u)\!:=\!u\tau(i,\alpha^{d},C^{\prime},P)$ (19)
of each rule
$r_{i}:=r_{i}^{\mathrm{P}}(v,P,u):=(v,w^{\mathrm{P}}(i,\alpha^{d},C^{\prime},P,u))$ (20)
for $i\in[1,m]$ is defined by a non-vanishing factor $u\in\mathbb{R}_{>0}$ and a pattern
$P:=\{(d_{1},v_{1}),\dots,(d_{n},v_{n})\}\in\mathcal{P}$ (21)
consisting of a set of $n\leq D$ direction-value pairs with $d_{i}\in[1,D]$ for $i\in[1,n]$, $d_{i}\neq d_{j}\,\forall\,i,j\in[1,n]\land i\neq j$, and $v_{i}\in A$ for $i\in[1,n]$, where
$\mathcal{P}:=\{P\,|\,P\in([1,D]\times A)^{n}\,\forall\,n\in[0,D]\}$ (22)
denotes the set of all possible patterns. A pattern $P$ is fulfilled, if
$\tau(i,\alpha^{d},C^{\prime},P):=\prod_{(d,v)\in P}\prod_{\begin{subarray}{c}(s,v^{\prime})\in C^{\prime}\\ \alpha^{d}_{i,s}=1\end{subarray}}\delta_{vv^{\prime}}\in[0,1]$ (23)
evaluates to one. Here, $\delta$ represents the Kronecker delta. The set of all possible pattern-based rulesets is denoted by $\mathcal{R}^{\mathrm{P}}\subset\mathcal{R}$.
Summarized, the weight function, eq. 19, yields $u$ if the pattern $P$ applies and zero otherwise. In line with expectations, only adjacent segments (i.e., $\alpha^{d}_{i,s}=1$) have an influence on eq. 16 according to eq. 23.
Furthermore, the rules from eq. 20 can be rewritten as
$r_{i}^{\mathrm{P}}(v,P,u)=(v,u,P)$ (24)
for $i\in[1,m]$ with a value $v\in A$, a factor $u\in\mathbb{R}_{>0}$, and a pattern $P\in\mathcal{P}$. For the pattern-based value selection, eq. 16 can be reformulated as
$p_{k}^{\mathrm{v,R}}(\nu|i,C^{\prime})$ $=p_{k}^{\mathrm{v,P}}(\nu|i,C^{\prime})$
$:=\sum_{\begin{subarray}{c}(v,u,P)\in R\\ v=\nu\end{subarray}}\gamma(u,i,C^{\prime})\tau(i,\alpha^{d},C^{\prime},P)$ (27)
with the coefficients
$\gamma(u,i,C^{\prime}):=\frac{u}{\overline{F}_{k}(i,C^{\prime})}\in[0,1]$ (28)
that correspond to the normalized weights of the rules.
In the context of this manuscript, one variant (that lies between rule- and pattern-based selection) is particularly important, namely the use of a functional factor instead of a constant factor for pattern-based rules of the form of eq. 24. Specifically, a factor
$u:=u(i,C):[1,N]\times\overline{\mathcal{C}}\rightarrow\mathbb{R}$ (29)
is considered that can depend both on the coordinate $i\in[1,N]$ and/or the already generated content $C\in\overline{\mathcal{C}}$, eq. 15. By mapping to zero, this functional factor can disable a pattern for certain coordinates or partial contents.
1.2.2 Identifier selection.
For the entropic segment identifier selection, the probability distribution from eq. 5 is defined as
$p_{k}^{\mathrm{s}}(i|C^{\prime})=p_{k}^{\mathrm{s,H}}(i|C^{\prime}):=\begin{cases}|I_{k}(C^{\prime})|^{-1}&\text{if }i\in I_{k}(C^{\prime})\\ 0&\text{otherwise}\end{cases}$ (30)
with the set of optimal solutions
$I_{k}(C^{\prime}):=\{i\,|\,i\in S^{-}(C^{\prime})\land H_{k}(i,C^{\prime})=h_{k}(C^{\prime})\}$ (31)
of the optimization task
$h_{k}(C^{\prime}):=\operatorname*{arg\,min}_{j\in S^{-}(C^{\prime})}H_{k}(j,C^{\prime}),$ (32)
where
$H_{k}(i,C^{\prime}):=-\sum_{v\in A}$ $p_{k}^{\mathrm{v,P}}(\nu=v|i,C^{\prime})$
×$\times\ln p_{k}^{\mathrm{v,P}}(\nu=v|i,C^{\prime})$ (33)
denotes the Shannon information entropy for $C^{\prime}\in\mathcal{C}(k-1)$, $k\in[1,N]$ That is, the segments with the smallest Shannon entropy with respect to the available value options are chosen.
1.3 CWFC Example: Checkerboard
As a simple example, consider the generation of images with black and white checkerboard patterns. The pixels of an image constitute the segments and the alphabet consists of only $W=2$ symbols standing for the two colors. As an adjacency configuration, the nearest-neighbor adjacency from fig. 1 is presumed.
A checkerboard image can be achieved with only two patterns of the form of eq. 21, namely $P_{1}:=\{(1,1),(2,1),(3,1),(4,1)\}$ and $P_{2}:=\{(1,2),(2,2),(3,2),(4,2)\}$. The corresponding ruleset, eq. 12, reads $R=\{r_{1}^{\mathrm{P}},r_{2}^{\mathrm{P}}\}$ and therefore contains the $m=2$ rules $r_{1}^{\mathrm{P}}(v=2,P=P_{1},u=1)$ and $r_{2}^{\mathrm{P}}(v=1,P=P_{1},u=1)$ based on eq. 20, see fig. 2.
Figure 2: Visualization of pattern-based rules. (a) Visualization scheme. The center tile corresponds to the target segment $i$ of the weight function, eq. 19, with a target value of $v$, whereas the adjacent tiles represent the required values $v_{1},\dots,v_{4}\in\{1,2\}$ of the four adjacent segments in the respective directions $d=1,\dots,d=4$. (b) Pattern-based rules $r_{1}^{\mathrm{P}}(v=2,P=P_{1},u=1)$ and $r_{2}^{\mathrm{P}}(v=1,P=P_{2},u=1)$, eq. 20, that can be used for the generation of images with black and white checkerboard patterns.
The iterative process of CWFC is sketched in fig. 3 for a $3\times 3$ image (represented by $N=9$ segments). Only two content instances can be generated, $C_{1}:=\{(1,1),\dots\}$ and $C_{2}:=\{(1,2),\dots\}$ with equal probability, eq. 10. The choice of the first value $v_{s(1)}$ already determines the choices of all other values.
Figure 3: CWFC for the generation of $3\times 3$ images with checkerboard patterns defined by the rules from fig. 2. In each iteration $k\in[1,9]$, three steps take place: (i) a segment identifier $s(k)$ is drawn, eq. 30, (ii) a corresponding value $v_{s(k)}$ is drawn, section 1.2.1, and (iii) the newly generated identifier-value pair is added as a segment to the content instance $C(k)$. For $s(k)$ and $v(k)$, the uniformly distributed set of possible choices are listed. The choices for $s(k)$ depend on the Shannon entropy $H_{k}$, section 1.2.2, of each undefined segment, as shown on the right. The choices for $v(k)$ depend on the fulfillment of the two rules $r_{1}^{\mathrm{P}}$ and $r_{2}^{\mathrm{P}}$, as listed.
2 QUANTUM METHODS
Based on the previously established classical concepts, a QWFC method and a HWFC method are proposed in the following. A software implementation is available online [9].
2.1 Quantum Wave Function Collapse (QWFC)
So far, the procedural generation concepts have been presented from a purely classical perspective. However, it seems natural that the probability distribution of the content instances $p(C)$, eq. 10, can also be realized with the help of a quantum circuit by utilizing the intrinsic randomness of quantum physics [10]. This is the very idea on which QWFC is based.
A single alphabet $A$ with $W$ symbols is presumed, eq. 11. Furthermore, two procedures control the content generation, namely segment identifier selection and value selection.
2.1.1 Identifier selection.
The Identifier selection is performed with a predefined order given by the vector $\sigma\in S_{N}$, which is the set of all permutations of $[1,N]$. Hence,
$p_{k}^{\mathrm{s}}(i|C^{\prime})=p_{k}^{\mathrm{s,o}}(i|C^{\prime}):=\delta_{i\sigma_{i}}.$ (34)
This approach leads to (partial) content instances $\mathcal{C}(k)=\mathcal{C}_{\sigma}(k)$, eq. 2, of the form
$\mathcal{C}_{\sigma}(k):=\{C\,|\,$ $C=\{(\sigma_{1},v_{\sigma_{1}}),\dots,(\sigma_{k},v_{\sigma_{k}})\}$
$\,\forall\,v_{\sigma_{l}}\in A\,\forall\,l\in[1,k]\}$ (35)
for $k\in[1,N]$ and $\mathcal{C}_{\sigma}(0):=\{\{\}\}$.
2.1.2 Value selection.
The value selection is pattern-based in analogy to section 1.2.1. In particular, the value for each segment is encoded by a set of qubits, which requires
$q:=\lceil\log_{2}W\rceil$ (36)
qubits for each segment $i\in[1,N]$. The group of $q$ qubits that represent the value for segment $i$ is denoted by
$\mathcal{Q}_{i}:=\{q_{i}^{1},\dots,q_{i}^{q}\},$ (37)
where $q_{i}^{j}$ references the $j$th qubit of this group for $j\in[1,q]$. The state of these qubits encodes the value $v_{i}$ of the segment $i$ in binary as the tensor product sequence of qubit states
$\bigotimes_{j=1}^{q}\ket{b_{j}}=\Ket{v_{i}-1=\sum_{j=1}^{q}b_{j}2^{j-1}}_{i}=\ket{v_{i}-1}_{i}\in\mathcal{H}_{i}$ (38)
with $b_{j}\in\mathbb{B}$, $v_{i}\in A$, and $i\in[1,N]$, where $\mathcal{H}_{i}$ denotes the joint Hilbert space of all qubit from $\mathcal{Q}_{i}$.
In total,
$Q:=Nq$ (39)
qubits are required to represent the value of all $N$ segments. The joint Hilbert space of all qubits from $\bigcup_{i=1}^{N}\mathcal{Q}_{i}$ is given by $\mathcal{H}:=\bigotimes_{i=1}^{N}\mathcal{H}_{i}$ and a state |$\ket{\Psi}\in\mathcal{H}$ can realize samples from a specific probability distribution of content instances
$p(C\!=\!\{(1,v_{1}),\dots,(N,v_{N})\})=|\bra{\Psi}\bigotimes_{i=1}^{N}\ket{v_{i}-1}_{i}|^{2}.$ (40)
based on a projective measurement. After these preliminary considerations, the remaining question is how such a state can be prepared constructively with quantum gates from a given pattern-based ruleset. This question will be answered in the following.
2.1.3 Quantum circuit.
Initially, all qubits are prepared in the ground state, i.e., |$\ket{\Psi}=\ket{0}:=\bigotimes_{i=1}^{N}\ket{0}_{i}$. Then, given a pattern-based ruleset $R\in\mathcal{R}^{\mathrm{P}}$, eq. 12, a series of operators is applied, one for each iteration $k\in[1,N]$ such that |$\ket{\Psi}:=U_{N}\cdots U_{1}\ket{0}$ represents the final state after $k$ iterations. In each iteration $k$, the state of the qubits $\mathcal{Q}_{\sigma_{k}}$ (that represent the value for segment $\sigma_{k}$) is prepared conditioned on the state of (a subset of) the qubits $\mathcal{Q}_{\sigma_{1}},\dots,\mathcal{Q}_{\sigma_{k-1}}$ (that represent the values of the already prepared segments $\sigma_{1},\dots,\sigma_{k-1}$), which leads to an entangled joint state. A sketch is shown in fig. 4 (a).
Figure 4: QWFC circuit with $\sigma=(1,\dots,N)$. (a) Circuit layout with an exemplarily alphabet $A$, eq. 11, with $W=16$ symbols, which requires a set $\mathcal{Q}_{i}$, eq. 37, of $q=4$ qubits, eq. 36, for each segment with $i\in[1,N]$. In each iteration $k\in[1,N]$, the operator $U_{k}$, eq. 41, prepares the qubits from $\mathcal{Q}_{k}$ in a superposition state |$\ket{C^{\prime}}_{k}$, eq. 45, entangled with (a subset of) the qubits from $\mathcal{Q}_{1},\dots,\mathcal{Q}_{k-1}$. All qubits are measured to obtain $p(C)$, eq. 40. (b) Gate decomposition of $U_{k}$ into the components of $U_{k}^{\mathrm{coin}}$, eq. 42, which consists of pairs $U_{k}^{\mathrm{ctrl}}(C^{\prime})\otimes U_{k}^{\mathrm{init}}(C^{\prime})$, eqs. 43 and 44. Each of these pairs represent a conditional loading of a probability distribution to realize |$\ket{C^{\prime}}_{k}$ (which is not explicitly shown). The combined control operators act on the qubits from $\mathcal{Q}_{k}^{+}$ eq. 47, whereas the initialization operators act on the qubits from $\mathcal{Q}_{k}$.
Specifically,
$U_{k}:=\left[U_{k}^{\mathrm{coin}}+U_{k}^{0}\right]\otimes\bigotimes_{l=k+1}^{N}\mathbbm{1}_{\sigma_{l}}\in\mathcal{H}$ (41)
for $k\in[1,N]$ contains $\mathbbm{1}_{i}\in\mathcal{H}_{i}$, the unit operator for $i\in[1,N]$ and two other operators. First, the conditional initialization operator
$U_{k}^{\mathrm{coin}}=\sum_{C^{\prime}\in\mathcal{C}_{\sigma}(k)}\left[U_{k}^{\mathrm{ctrl}}(C^{\prime})\otimes U_{k}^{\mathrm{init}}(C^{\prime})\right],$ (42)
where $\mathcal{C}_{\sigma}(k)$ represents the set of possible partial content instances from section 2.1.1,
$U_{k}^{\mathrm{ctrl}}(C^{\prime}):=\bigotimes_{(s,v)\in L_{k}(C^{\prime})}\ket{v-1}\bra{v-1}_{s}\in\bigotimes_{l=1}^{k-1}\mathcal{H}_{\sigma_{l}}$ (43)
denotes the control operator, and
$U_{k}^{\mathrm{init}}(C^{\prime}):=\ket{C^{\prime}}\bra{0}_{\sigma_{k}}\in\mathcal{H}_{\sigma_{k}}$ (44)
the initialization operator with
|$\ket{C^{\prime}}_{\sigma_{k}}:=\sum_{\nu\in A}\sqrt{p_{k}^{\mathrm{v,P}}(\nu|\sigma_{k},C^{\prime})}\ket{\nu-1}_{\sigma_{k}}\in\mathcal{H}_{\sigma_{k}},$ (45)
which is based on the pattern-based probability distribution of values $p_{k}^{\mathrm{v,P}}(\nu|i,C^{\prime})$ from section 1.2.1. The summation in eq. 43 runs over all adjacent segments
$L_{k}(C^{\prime}):=\{(s,v)\,|\,$ $(s,v)\in C^{\prime}\,\land\,\exists\,P\in R_{P}$
$\exists\,d\in P_{d}:\alpha^{d}_{\sigma_{k}s}=1\}\subseteq C^{\prime},$ (46)
where $R_{P}:=\{P_{1},\dots,P_{m}\}$ stands for the set of patterns in the pattern-based ruleset $R\in\mathcal{R}^{\mathrm{P}}$ with $P_{i}\in\mathcal{P}$ for $i\in[1,m]$ and $P_{d}:=\{d_{1},\dots,d_{n}\}$ stands for the set of directions in the pattern $P\in\mathcal{P}$, eq. 21, with $d_{j}\in[1,D]$ for $j\in[1,n]$. In particular, as discussed above for the classic case, choosing $C^{\prime}$ instead of $L_{k}(C^{\prime})$ in the summation has no effect on $p_{k}^{\mathrm{v,P}}(\nu|i,C^{\prime})$ and thus |$\ket{C^{\prime}}_{\sigma_{k}}$.
In iteration $k$, only the qubits in $\mathcal{Q}_{k-1}^{+}$ with
$\mathcal{Q}_{k}^{+}:=\{q\,|\,q=q_{k}^{i}\in\mathcal{Q}_{\sigma_{k}}\,\forall\,i\in[1,q]\,\land\,\sigma_{k}\in L_{k}^{+}\}$ (47)
are affected by $U^{\mathrm{ctrl}}(C^{\prime})$ for some $C^{\prime}\in\mathcal{C}_{\sigma}(k)$, where
$L_{k}^{+}:=\{s\,|\,\exists\,C^{\prime}\in\mathcal{C}_{\sigma}(k)\,:\,s\in S^{+}(L_{k}(C^{\prime}))\}$ (48)
denotes the affected segments for $k=[1,N]$. Here, $\mathcal{Q}_{k}^{+}\subseteq\{\mathcal{Q}_{1},\dots,\mathcal{Q}_{k}\}$ and $L_{k}^{+}\subseteq\{\sigma_{1},\dots,\sigma_{k}\}$, respectively.
The second operator in eq. 41 is the neutral operator
$U_{k}^{0}=\sum_{C^{\prime}\in\mathcal{C}_{\sigma}(k)}\left[U_{k}^{\mathrm{ctrl}}(C^{\prime})\otimes U_{k}^{\mathrm{init},0}+U_{k}^{\mathrm{ctrl},0}(C^{\prime})\right].$ (49)
Initially, all qubits $\mathcal{Q}_{k}$ are in the ground state. Therefore,
$U_{k}^{\mathrm{init},0}:=\sum_{v\in[2,W]}\ket{v-1}\bra{v-1}_{\sigma_{k}}\in\mathcal{H}_{\sigma_{k}}$ (50)
vanishes. Furthermore,
$U_{k}^{\mathrm{ctrl},0}(C^{\prime})\!:=\!\left[\bigotimes_{l=1}^{k-1}\mathbbm{1}_{\sigma_{l}}-U_{k}^{\mathrm{ctrl}}(C^{\prime})\right]\otimes\mathbbm{1}_{\sigma_{k}}\!\in\bigotimes_{l=1}^{k}\mathcal{H}_{\sigma_{l}}$ (51)
does not change the state by construction.
The operator $U_{k}$, eq. 41, can be straightforwardly realized with quantum gates for a given $k\in[1,N]$. To this end, only the conditional initialization operator $U_{k}^{\mathrm{coin}}$, eq. 42, has to be considered because all other parts of $U_{k}$ have no effect on the resulting state. The conditional initialization operator consist of a sequence of joint pairs of control-initialization operators $U_{k}^{\mathrm{ctrl}}(C^{\prime})\otimes U_{k}^{\mathrm{init}}(C^{\prime})$, one for each $C^{\prime}\in\mathcal{C}_{\sigma}(k)$. Each of these joint pairs represent a conditional loading of a probability distribution to realize |$\ket{C^{\prime}}_{\sigma_{k}}$, which is a well-known task that may, however, require an exponential number of gates [11, 12, 13]. A sketch is shown in fig. 4 (b).
For the definition of QWFC, a noise-free, ideal quantum device is presumed, that is able to run the emerging circuits without any errors. However, the currently available noisy intermediate-scale quantum (NISQ) hardware exhibits significant noise and uncertainty [14]. As a consequence, the content sampled from eq. 40 can also be expected to violate the prescribed patterns. This will be demonstrated further below for several PCG use cases.
The proposed conditions S1, S2, V1, and V2 are in fact not mandatory in the original implementation of CWFC, where a violation leads to a restart of the algorithm. This can also be considered in QWFC by adding an additional “conflict detection qubit” that stores this information, an extension that is not further discussed here.
2.2 QWFC Example: Checkerboard
In the following, the above example for the generation of $3\times 3$ checkerboard images is considered for QWFC, where the predefined segment order $\sigma=(1,2,3,6,5,4,7,8,9)$ is used for eq. 34 as visualized in fig. 5 (a). The resulting CWFC circuit is shown in fig. 5 (b) and the resulting probability distribution of content instances $p(C)$, eq. 40, in fig. 5 (c). Given a noise-free quantum device, a measurement of the circuit corresponds to drawing a sample from this distribution.
Figure 5: QWFC for the generation of $3\times 3$ images with black and white checkerboard patterns in analogy to fig. 3. (a) Predefined segment order $\sigma=(1,2,3,6,5,4,7,8,9)$ for eq. 34. (b) Circuit layout with the same symbols as in fig. 4. (c) Generated content instances $C_{1}$ and $C_{2}$ with $p(C_{1})=p(C_{2})=\frac{1}{2}$, eq. 40.
2.3 Hybrid Quantum-Classical WFC (HWFC)
The required number of qubits for the proposed QWFC method increases linearly with the number of segments and logarithmically with the number of symbols in the alphabet as shown in eqs. 36 and 39. At the same time, the respective circuits may become exponentially deeper because of the costly conditional preparation of probability distributions. The reliable execution of such big circuits on NISQ hardware can be very challenging or even impossible.
[…]
Finally, the distribution of the joint content instances is given by
$p\left(C=\bigcup_{h=1}^{H}C^{h}\right)=\prod_{h=1}^{H}p^{h}(C^{h}|C^{1},\dots,C^{h-1}),$ (52)
which corresponds to performing a QWFC on the complete content, cf. fig. 6. With the proposed HWFC method, the size of the QWFC circuits can be reduced at the price of an additional outer iteration loop and the inclusion of classical correlations in the conditional QWFC circuit generation.
Figure 6: HWFC with $H=3$ partitions of $N=9$ segments. For each partition, a conditional QWFC is performed and the resulting content instances are combined.
3 DEMONSTRATION
In the present section, five simple content creation use cases are presented to demonstrate the proposed QWFC and HWFC methods. For the realization of these examples, both quantum computing simulators and actual IBM Quantum devices were used, which were accessed via the IBM Quantum cloud services [15] during December 2023.
3.1 Checkerboard Revisited
As a first use case, the previously presented checkerboard example is recalled, cf. figs. 2 and 5. Again, $3\times 3$ instances are considered, for which a comparison between QWFC and HWFC is performed on the IBM Quantum device ibm_kyoto (Eagle r3 processor version 1.2.4, 127 qubits). For QWFC, 10 0001000010\,00010 000 shots are measured, each yielding a content instance. For HWFC, $H=3$ equal-sized partitions are used in analogy to fig. 6 and the method is repeated $131$ times, each time yielding a content instance. The results are shown in fig. 7.
[…]
In total, there are $5122^{9}=$ 512 possible content instances $C\in\mathcal{C}$, eq. 1, but only two of them, $C_{1}$ and $C_{2}$, are valid. The remaining $510$ content instances are invalid because they violate the patterns. In fig. 7 (a), QWFC is able to generate the two valid instances, but at the same time also produces a lot of invalid instances due to hardware imperfections. From fig. 7 (b) it can be seen that the results from HWFC are much closer to the idealized solution that would contain only $C_{1}$ and $C_{2}$ with equal probability. It is a consequence of the reduced circuit size from this method that invalid solutions occur with much less frequency than for QWFC.
3.2 Pipes
The second use case is to create an image from tiles of seven different pipe (or line) sections as well as one blank tile. That is, the alphabet consists of $W=8$ symbols and each tile in the image constitutes a segment. The alphabet is shown in fig. 8 (a). As adjacency configuration, the nearest-neighbor adjacency from fig. 1 is presumed with $D=4$. The pattern-based ruleset is chosen in such a way that neighboring tiles have to form a connecting network of pipes. Three example rules for this purpose are shown in fig. 8 (b). In total, $m=2048$ rules are required to achieve this goal.
Figure 8: Pipe generation. The generated image is required to show a network of connected pipes. (a) Alphabet consisting of $W=8$ symbols, each of which represents a tile with a pipe section and an empty tile. (b) Example of three pattern-based rules $r_{1}^{\mathrm{p}}$, $r_{2}^{\mathrm{p}}$, and $r_{3}^{\mathrm{p}}$, cf. fig. 2. (c) Example $10\times 4$ tile images that were generated using HWFC with $H=10$ on the IBM Quantum device ibm_kolkata and on simulator as well as CWFC.
Exemplarily, images consisting of $10\times 4$ tiles ($N=40$) are considered, for which a comparison between HWFC on the IBM Quantum device ibm_kolkata (Falcon r5.11 processor version 1.14.8, 27 qubits), HWFC on a simulator and CWFC is performed. For HWFC, $H=10$ equal-sized partitions are used, each corresponding to one column of the image. A few resulting instances are shown in fig. 8 (c).
Based on a visual comparison, both CWFC and HWFC on the simulator yield similar results, as expected. However, the results from ibm_kolkata clearly contain a lot of invalid patterns as a result of the imperfect hardware. This demonstrates the limitations of the proposed method on NISQ devices.
3.3 Hexagon Map
The third use case addresses images based on hexagonal tiles that can be interpreted as a map consisting of different terrains. For this purpose, an alphabet of $W=4$ symbols is chosen, each corresponding to a unicolored tile (blue, yellow, green, and gray) and each tile constitutes a segment. As adjacency configuration, a nearest-neighbor adjacency is presumed ($D=6$). The prescribed pattern-based ruleset only allows connections of the form blue-yellow-green-gray, which can be fulfilled with $m=1586$ rules, see fig. 9 (a). To promote a higher occurrence of blue tiles, the factors in eq. 24 are chosen as $u=5$ for blue tiles and $u=1$ otherwise.
Figure 9: Map generation. Only connections of the form blue-yellow-green-gray are allowed. (a) Example of three pattern-based rules $r_{1}^{\mathrm{p}}$, $r_{2}^{\mathrm{p}}$, and $r_{3}^{\mathrm{p}}$. (b) Distribution of content instances $p(C)$ for $N=7$ hexagons using QWFC on a simulator (sorted by descending probability) and CWFC (with the same ordering). The inset plot shows the predefined segment order $\sigma$ and the five most probable instances for QWFC. The asymmetry is a consequence of the choice of $\sigma$. (c) Example images that were generated from $N=127$ hexagons using HWFC with $H=20$ on a simulator.
As a first example, images consisting of $N=7$ tiles are considered, for which QWFC on a simulator and CWFC are performed. The resulting distributions of content instances $p(C)$ in fig. 9 (b) is exact for QWFC, eq. 40, and based on 10 0001000010\,00010 000 samples for CWFC, eq. 10. Without surprise, both distributions are significantly different. The second example considers images from $N=127$ tiles, for which HWFC with $H=20$ is performed on a simulator. Example instances are shown in fig. 9 (c).
3.4 Platformer
The fourth use case is motivated by computer game level design. The goal is to create a platformer-type level from a set of eight tiles [16], each corresponding to a different game element (ground, grass, mushroom, block, air, and a tree that is composed of three tiles). Consequently, the alphabet consists of $W=8$ symbols, see fig. 10 (a). As adjacency configuration, the two tiles above and below are taken into account as neighbors ($D=2$). A pattern-based ruleset is prescribed that meets five requirements: (i) The bottom row consists of ground tiles and ground tiles can only be placed on top of each other, (ii) only a grass tile or a mushroom tile can be placed above a ground tile, (iii) the tree tiles must be placed in order with the lowest tile above a grass tile, (iv) air tiles can only be placed above grass, mushroom or other air tiles, (v) block tiles can only be placed between two air tiles. These requirements can be fulfilled with $m=15$ rules, see fig. 10 (b). The first requirement can be resolved with a functional factor, eq. 29, that vanishes for all non-ground tiles in the bottom row. To reduce the occurrence of block tiles, the corresponding factor in eq. 24 is chosen as $u=$0.1$$ for $r_{6}^{\mathrm{p}}$. All other factors are set to $u=1$. The segment order $\sigma$ is chosen such that the levels are built up from bottom to top.
Figure 10: Level generation. (a) Alphabet consisting of eight symbols, each of which represents a tile [16]. (b) Pattern-based ruleset consisting of $m=15$ rules. (c) Example $10\times 10$ tile images that were generated using HWFC with $H=20$ on a simulator.
Exemplarily, levels with $10\times 10$ tiles are considered, for which HWFC with $H=20$ equal-sized partitions (each corresponding to a half row of the level) is performed on a simulator. Four resulting instances are shown in fig. 10 (c).
3.5 Voxel Skyline
The final use case considers the creation of a 3d voxel graphic with a binary alphabet of $W=2$ symbols, which represent the presence or absence of a voxel. Each voxel constitutes a segment, as adjacency configuration the voxel above and below are taken into account ($D=2$). The chosen pattern-based ruleset with $m=4$ ensures that voxels are built from the ground up, see fig. 11 (a). This leads to a skyline-like structure.
Figure 11: Skyline use case. (a) Pattern-based ruleset consisting of $m=4$ rules in an analogous representation as in the previous figures. (b) Example images from a $4\times 4\times 4$ voxel grid that were generated using QWFC on the IBM Quantum device ibm_osaka and a simulator. (c) Example images from a $10\times 10\times 5$ voxel grid that were generated using QWFC on a simulator.
Exemplarily, two cases are considered: First, images from a $4\times 4\times 4$ voxel grid, for which QWFC is performed both on a the IBM Quantum device ibm_osaka (Eagle r3 processor version 1.0.3, 127 qubits) and a simulator, see fig. 11 (b). An error violating $r_{3}^{\mathrm{p}}$ can be seen in the bottom left image that results from hardware imperfections. Second, images from a $10\times 10\times 5$ voxel grid, for which QWFC is performed on a simulator, see fig. 11 (c).
4 CONCLUSIONS
In this work, QWFC was proposed as a method to realize WFC on a quantum computer for PCG. The idea is to construct a quantum circuit from which valid content instance (that fulfill the prescribed patterns) can be sampled via measurements. In other words, the quantum circuit acts as a special-purpose random generator for content of a desired form that makes use of the intrinsic randomness of quantum physics. This method is not without its challenges. The corresponding quantum circuit becomes larger from both the content and alphabet size and at the same time deeper from the correlations within the content instance distribution, which makes it difficult to evaluate most use cases on current hardware. For this reason, the quantum-classical hybrid method HWFC was proposed that decomposes the problem into smaller sub-problems which can then be executed on the quantum computer with a higher chance of success.
The experimental results have shown the limitations of QWFC and HWFC, but have also proven that at least simple examples can already be implemented on today’s quantum hardware. Furthermore, all experiments were carried out without quantum error correction [17], which can be used to possibly improve the results. The biggest bottleneck of the circuit complexity is the conditional preparation of probability distributions. Investigating possible improvements in this area with regard to specific PCG use cases or alternative preparation approaches (e.g., qGANs [18]) could therefore be a potential direction for future research.
The conceptual advantage of QWFC is that a circuit, once prepared, represents the distribution of content instances. This means that new instances can be generated by simple measurements without any additional algorithmic effort. A shortcoming of QWFC is that only a fixed order of segments is considered instead of a dynamic (e.g., entropy-based) order that depends on the previously selected segments. The realization of such a generalization of the method could also be a point of origin for further studies.
CWFC is a strong method with many areas of application. In this paper, only visual content creation use cases were presented, but QWFC can also be used for non-visual contents like text [9] in analogy to CWFC [19].
In conclusion, the paper aims to provide an initial approach to QWFC that demonstrates its feasibility but leaves room for improvement and raises the question of useful application areas given the limits of NISQ devices.
5 ACKNOWLEDGMENTS
Parts of this research have been funded by the Ministry of Science and Health of the State of Rhineland-Palatinate (Germany) as part of the project AnQuC.

### v2

@@
-Quantum Wave Function Collapse for Procedural Content Generation
+angle=0, scale=1, contents= This article has been accepted for publication in IEEE Computer Graphics and Applications. This is the author’s version which has not been fully edited and
 Abstract
@@
 This quantum wave function collapse algorithm is based on the idea that a quantum circuit can be prepared in such a way that it acts as a special-purpose random generator for content of a desired form.
-The proposed method is presented theoretically and investigated experimentally on simulators and actual IBM Quantum devices.
+The proposed method is presented theoretically and investigated experimentally on simulators and IBM Quantum devices.
+[…]
 Wave function collapse (WFC) is a powerful tool for procedural content generation (PCG) that is for example used in video games, as it can save significant development time by automating the creation of diverse and complicated game elements.
@@
 Content in this context can therefore have a broad range of meanings: Images, 3D models, game levels, text, sound or a combination of these, to name just a few examples.
-Originally proposed in Ref.
-[1], the WFC algorithm is a a non-backtracking, greedy constraint solving method [2, 3] that is able to generate complex patterns based on a set of input samples.
+The WFC algorithm [1] is a a non-backtracking, greedy constraint solving method [2, 3] that is able to generate complex patterns based on a set of input samples.
 It is known for its ability to create diverse and complex outputs that resemble the input samples while exhibiting novel combinations and variations.
@@
 Despite its name, WFC is a purely classical algorithm.
-But even if WFC has nothing to do with quantum physics beyond the terminology, the question naturally arises as to whether quantum computers can be used to execute a genuine quantum version of the algorithm.
-The investigation of this question and a first draft for a corresponding algorithm are the subject of this manuscript.
-For this purpose, the terms classical wave function collapse (CWFC), quantum wave function collapse (QWFC), and hybrid quantum-classical wave function collapse (HWFC) are used In the following to clearly distinguish between classical, quantum and hybrid quantum-classical WFC.
-Leveraging quantum computers for PCG has been considered before, mainly focusing on the quantum generalization of a blurring process [5, 6] and map generation using a quantum-enhanced decision making process [7].
-Related projects have been carried out in connection with the development of quantum-inspired games [8].
-The contribution of this paper is threefold:
-  • A probabilistic formulation of the simple tiled model of CWFC is presented as a foundation for a quantum version.
-  • A QWFC method and a HWFC method are proposed.
-  • The proposed methods are tested on simulators and actual IBM Quantum devices.
-The remaining manuscript is structured as follows.
-In the next section, a formal description of CWFC is provided in an appropriate probabilistic form, which enables the development of a QWFC method and a HWFC method in the subsequent section.
-In the section after, the proposed methods are demonstrated in practice.
-Finally, the paper ends with a conclusion and an outlook.
-1 CLASSICAL METHODS
-In this section, the formal description for a probabilistic iterative PCG (PIPCG) is provided and, subsequently, it is shown how it can be reduced to the common CWFC approach.
-1.1 PIPCG
+But even if WFC has nothing to do with quantum physics beyond the terminology, the question arises as to whether quantum computers can still be used to execute a genuine quantum version of the algorithm.
+The motivation for leveraging quantum computers for PCG, which is called quantum procedural content generation (QPCG) in the following, lies in their potential to introduce new levels of complexity, creativity, and efficiency.
+Because quantum physics exhibits intrinsic randomness [5], quantum random number generators (QRNGs) can be viewed as sources of true randomness [6], making them a natural choice for generating randomized content.
+In addition, quantum superposition allows the random generation of complex patterns and variations that classical algorithms might find challenging to generate efficiently.
+Currently available noisy intermediate-scale quantum (NISQ) hardware has very limited capabilities and is subject to significant noise and uncertainty [7].
+While these kind of disruptive effects are typically to be avoided, they can also lead to novel and unexpected results for the generated content, potentially fostering creativity.
+In this sense, despite the technological limitations, it seems promising to use quantum computers already today for simple PCG tasks and to study the results.
+For more complex tasks, hybrid quantum-classical algorithms are a promising strategy to overcome the limitations, for example by integrating quantum subroutines into classical PCG methods to increase variability and complexity.
+QPCG is therefore a promising approach, currently mainly for research, but in the future possibly also for practical applications.
+However, a general discussion of QPCG is beyond the scope of this paper.
+Instead, the focus will specifically be on the WFC algorithm and how it can be modified to enable the use of gate-based quantum computers [8].
+This paper has four main contributions:
+  • A probabilistic formulation of the simple tiled model of WFC is presented as a foundation for a quantum version.
+We denote this formulation as classical wave function collapse (CWFC) to clearly distinguish between classical and non-classical approaches.
+  • The quantum wave function collapse (QWFC) method is proposed, which founds on the encoding of a probability distribution in a quantum circuit.
+  • The hybrid quantum-classical wave function collapse (HWFC) method is proposed to take into account the limitations of NISQ hardware.
+  • The proposed methods are tested on simulators and IBM Quantum devices.
+The remainder of the manuscript is organized as follows.
+After a brief summary of related work, a formal description of CWFC is given in an appropriate probabilistic form, which subsequently allows the development of QWFC and HWFC.
+The proposed methods are then demonstrated in practice.
+Finally, the paper ends with a conclusion.
+RELATED WORK
+There are many approaches to PCG, such as machine learning [9], evolutionary algorithms [10] and search-based methods [11].
+WFC was originally proposed by Gumin [1] as an example of the latter.
+Subsequent research has extended the basic WFC algorithm to address its limitations and expand its capabilities [2, 3].
+For example, by eliminating the need for two-dimensional grids [12], adding design constraints [13], or allowing interactive user control over the results [14, 15].
+Better scaling and runtime can be achieved by a hierarchical approach [16], a nested approach [4] or by using bitwise operations [17].
+In contrast to classical PCG, QPCG is much less common in the literature.
+Previous work on QPCG was mainly focusing on the quantum generalization of a blurring process [18, 19] and map generation using a quantum-enhanced decision making process [20].
+Moreover, related projects have been carried out in connection with the development of quantum-inspired games [21].
+CLASSICAL METHODS
+This section provides a formal framework for the following methods, referred to as probabilistic iterative PCG (PIPCG).
+Subsequently, it is shown how the common CWFC is a special case of PIPCG.
+Only the main features of PIPCG and CWFC are summarized, the technical details can be found in the appendix.
+Probabilistic Iterative PCG (PIPCG)
 To begin with, it is first necessary to formally define the generated content at a sufficiently abstract level.
@@
 $C:=\{(1,v_{1}),\dots,(N,v_{N})\}\in\mathcal{C},$ (1)
-where each segment $(i,v_{i})$ is defined by a unique identifier $i\in[1,N]$ and a value $v_{i}\in A_{i}$ from an alphabet $A_{i}\subset\mathbb{N}$.
-Here and in the following, the notation $[i,j]:=\{i,\dots,j\}\,\forall\,i,j\in\mathbb{N}$ is used.
-The set of all possible content instances $\mathcal{C}$ is determined by the content alphabets $\mathcal{A}:=\{A_{1},\dots,A_{N}\}$ for a given size $N$ with $|\mathcal{C}|=\prod_{i=1}^{N}A_{i}$.
+where each segment $(i,v_{i})$ is defined by a unique identifier $i\in[1,N]$ and a value $v_{i}\in A$ from an alphabet $A:=[1,W]$ consisting of $W$ symbols with the notation $[i,j]:=\{i,\dots,j\}$.
+The set of all possible content instances $\mathcal{C}$ contains $W^{N}$ elements.
 For example, if the content represents an image, its individual pixels can constitute the segments.
 Each pixel can attain one color from the available color space of the image (as its alphabet).
-1.1.1 Iterative process.
+Iterative Process.
 The procedural generation of content instance $C\in\mathcal{C}$ is realized through an iterative process, where each iteration $k\in[1,N]$ consists of three steps:
   1. 1.
-    Randomly choose a new segment identifier $s(k)\in[1,N]$ from a set of available identifiers.
+    Randomly select a new segment identifier $s(k)\in[1,N]$ from a set of available identifiers with a user-defined probability distribution.
   2. 2.
-    Randomly choose a corresponding value $v_{s(k)}\in A_{s(k)}$ from a set of available values.
+    Randomly select a corresponding value $v_{s(k)}\in A$ from a set of available values with a user-defined probability distribution.
   3. 3.
     Add the chosen segment $(s(k),v_{s(k)})$ to the content instance.
-By repeating these steps, the content instance is assembled in the sense of $C(0)\overset{k=1}{\rightarrow}C(1)\overset{k=2}{\rightarrow}\cdots\overset{k=N}{\rightarrow}C(N)$, where
-$C(k):=\{(s(1),v_{s(1)}),\dots,(s(k),v_{s(k)})\}\in\mathcal{C}(k),$ (2)
-represents the generated partial content instance at the end of iteration $k$ with $|C(k)|=k$, $C(k)=C(k-1)\cup\{(s(k),v_{s(k)})\}$, $s(i)\neq s(j)\,\forall\,i,j\in[1,k]\land i\neq j$.
-Here, $\mathcal{C}(k)$ denotes all possible partial content instances at the end of iteration $k$.
-Thus, $|\mathcal{C}(k)|=\sum_{I\in J(k)}\prod_{i\in I}|A_{i}|$, where $J(k):=\{I\,|\,I\in[1,N]\land|I|=k\}$ with $|J(k)|={N\choose k}$.
-It is convenient to define $C(0):=\{\}$ and $\mathcal{C}(0):=\{C(0)\}$.
-In addition,
-$S^{+}(C(l)):=\{s(1),\dots,s(l)\}$ (3)
-denotes the set of segment indices (dropping the values) that are contained in the partial content instance $C(l)$ with $|S^{+}(C(l))|=l$ and
-$S^{-}(C(l)):=[1,N]\setminus S^{+}(C(l))$ (4)
-denotes the remaining segment indices that are not contained in $C(l)$ with $|S^{-}(C(l))|=N-l$ for all $C(l)\in\mathcal{C}(l)$ and for all $l\in[0,N]$.
-After $N$ iterations, the newly generated content instance is complete, i.e., $C(N)\equiv C$ and $\mathcal{C}(N)\equiv\mathcal{C}$.
-In the following, the three steps within each iteration $k$ are explained in more detail.
-1.1.2 First step.
-For the first step, an identifier is selected by drawing a sample $s(k)$ from the random variable
-$\boldsymbol{s}(k)\sim p^{\mathrm{s}}_{k}(i=s(k)|C^{\prime}=C(k-1)).$ (5)
-The probability distribution $p_{k}^{\mathrm{s}}(i|C^{\prime})$ with support $i\in[1,N]$ and $C^{\prime}\in\mathcal{C}(k-1)$ is a user-defined parameter of the procedure that determines how new segment indices $l$ are selected based on the partial content instance $C^{\prime}=C(k-1)$ that has already been generated up to the previous iteration $k-1$.
-The dependence on all previously chosen segments makes the process generally non-Markovian.
-It is required that two conditions hold true for the probability distribution $p_{k}^{\mathrm{s}}(i|C^{\prime})$:
-  1. S1
-    No identifier can be chosen twice, i.e., $p_{l}^{\mathrm{s}}(i=s|C^{\prime}=C(l-1))=0\,\forall\,s\in S^{+}(C(l-1))\,\forall\,C(l-1)\in\mathcal{C}(l-1)\,\forall\,l\in[1,N]$.
-  2. S2
-    At least one valid identifier has to be available, i.e., $\forall\,l\in[1,N]\,\forall\,C(l-1)\in\mathcal{C}(l-1)\,\exists\,s\in S^{-}(C(l-1)):p_{l}^{\mathrm{s}}(i=s|C^{\prime}=C(l-1))>0$.
-The segment identifier selection mechanism is one of the main ingredients of CWFC.
-Here, it is presented in very generic way.
-The common entropic approach will be explained further below along with possible variants that are more suitable for a quantum approach.
-1.1.3 Second step.
-After the identifier $s(k)$ has been chosen, the second step is to select a corresponding value $v_{s(k)}$ for the segment. For this purpose, a sample $v_{s(k)}$ is drawn from the random variable
-$\boldsymbol{v}_{s(k)}\sim p^{\mathrm{v}}_{k}(\nu=v_{s(k)}|i=s(k),C^{\prime}=C(k-1)).$ (6)
-The probability distribution $p^{\mathrm{v}}_{k}(\nu|i,C^{\prime})$ with support $\nu\in\cup_{A\in\mathcal{A}}A$, $i\in[1,N]$ and $C^{\prime}\in\mathcal{C}(k-1)$ is a user-defined parameter of the procedure that determines how new values are selected based on the selected identifier $i=s(k)$ of the current iteration and the partial content instance $C^{\prime}=C(k-1)$ that has already been generated up to the previous iteration $k-1$.
-It is required that two conditions hold true for the probability distribution $p^{\mathrm{v}}_{k}(\nu|i,C^{\prime})$:
-  1. V1
-    Only values from the corresponding alphabet are available, i.e., $p^{\mathrm{v}}_{l}(\nu=v|i=s,C^{\prime}=C(l))=0\,\forall\,v\not\in A_{s}\,\forall\,s\in[1,N]\,\forall\,C(l)\in\mathcal{C}(l-1)\,\forall\,l\in[1,N]$.
-  2. V2
-    At least one value from the corresponding alphabet has to be available, i.e., $\forall\,l\in[1,N]\,\forall\,C(l-1)\in\mathcal{C}(l-1)\forall s\in S^{-}(C(l-1))\exists v\in A_{s}:p_{l}^{\mathrm{v}}(\nu=v|i=s,C^{\prime}=C(l))>0$.
-As explained further below, it is also possible to specify $p^{\mathrm{v}}_{k}(\nu|i,C^{\prime})$ implicitly based on a set of rules, which is the default approach for CWFC and much more convenient for practical purposes than an explicit definition.
-To simplify the notation a joint distribution for the segment identifier and value selection $p_{k}(i,\nu|C^{\prime}):=p_{k}^{\mathrm{s}}(i|C^{\prime})p_{k}^{\mathrm{v}}(\nu|i,C^{\prime})$ can be defined.
-1.1.4 Third step.
-In the third step, the selected tuple $(s(k),v_{s(k)})$ that constitutes the newly generated segment is added to the already generated segments of the content instance, i.e., $C(k)=C(k-1)\cup\{(s(k),v_{s(k)})\}$.
-If $k<N$, the process is repeated and another segment is generated for the next iteration $k+1$.
-If $k=N$, the PIPCG is finished.
-1.1.5 Instance distribution.
-An essential concept of PCG is, that only a limited subset of all possible content instances $C\in\mathcal{C}$ is generated according to the designer’s aesthetic preferences, a given rule set, or some other underlying logic.
-Otherwise, the instances could be drawn directly from $\mathcal{C}$ with much less effort.
-For the presented formalism, the probability to generate a partial content instance $C(k)$ in the form of eq.
-2 is given by
-$p(C(k)):=\sum_{\sigma\in S_{k}(C(k))}\prod_{l=1}^{k}q_{l}^{k}(\sigma,C(k))$ (7)
-with the abbreviation
-$q_{l}^{k}(\sigma,C(k)):=p_{l}($ $i=\sigma_{l},\nu=v_{\sigma_{l}}|$
-$C^{\prime}\!=\!\{(\sigma_{1},v_{\sigma_{1}}),\!\dots,\!(\sigma_{l-1},v_{\sigma_{l-1}})\}).$ (8)
-Here, $S_{k}(C(k))$ denotes the set of all permutations of $S^{+}(C(k))$, eq.
-3, with $|S_{k}(C(k))|=k!$ and $\sigma_{l}$ represents the $l$th value of a permutation $\sigma\in S_{k}(C(k))$ for $l\in[1,k]$.
-The notation in section 1.1.5 is to be understood in such a way that $C^{\prime}=\{\}$ for $l=1$.
-Moreover, $v_{\sigma_{m}}$ in section 1.1.5 represents the value of the segment with identifier $\sigma_{m}$ from the content instance $C(k)$ for $m\in[1,l]$.
-By definition, the partial content instance $C(k)$ can only be generated if $p(C(k))>0$.
-In other words, the support
-$\underline{\mathcal{C}}(k):=\{C(k)\,|\,C(k)\in\mathcal{C}(k)\,\land\,p(C(k))>0\}$ (9)
-of $p(C(k))$ contains all partial content instances that can be generated for the given $k$.
-For $k=N$, consider a complete content instance $C$ in the form of eq. 1 with values $v_{1},\dots,v_{N}$. In this case, eq. 7 reads
-$p(C)=p(v_{1},\dots,v_{N}):=\sum_{\sigma\in S_{N}}\prod_{l=1}^{N}q_{l}^{N}(\sigma,C),$ (10)
-where $S_{N}:=S_{N}(C)$ is the set of all permutations of $[1,N]$.
-In fact, performing the presented PIPCG procedure corresponds to drawing a sample from this distribution such that the content instance can be considered as a random variable $\boldsymbol{C}\sim p(C)$.
-Consequently, $\underline{\mathcal{C}}(N)=\underline{\mathcal{C}}:=\{C\,|\,C\in\mathcal{C}\,\land\,p(C)>0\}$ denotes the set of all content instances that can be generated.
-1.2 Classical Wave Function Collapse (CWFC)
-CWFC is a special case of the previously described PIPCG.
-In this manuscript, only a simplified form of the simple tiled model of CWFC is considered, as it is sufficient to capture the key components of the approach.
-A detailed description of the original CWFC implementation can be found in, e.g., Refs.
-[2, 3].
-For the common CWFC approach, a single alphabet is used for all segments, i.e., $A_{i}=A\,\forall\,i\in[1,N]$. Given $W$ symbols, one can presume without loss of generality that
-$A:=[1,W].$ (11)
-Furthermore, the segment identifier selection is performed with an entropic approach, whereas the value selection is pattern-based.
-These two concepts will be explained in more detail in the following.
-1.2.1 Value selection.
+By repeating these steps, the content instance is assembled in the sense of $C(0)\overset{k=1}{\rightarrow}C(1)\overset{k=2}{\rightarrow}\cdots\overset{k=N}{\rightarrow}C(N)\equiv C$, where $C(k):=\{(s(1),v_{s(1)}),\dots,(s(k),v_{s(k)})\}$ represents the generated partial content instance at the end of iteration $k$.
+After $N$ iterations, the newly generated content instance is complete.
+Probabilistic Generation.
+An essential concept of PCG is that only a limited subset of all possible content instances $C\in\mathcal{C}$ is generated according to the designer’s aesthetic preferences, a given set of rules, or some other underlying logic; otherwise, the instances could be drawn directly from $\mathcal{C}$ with much less effort.
+For PIPCG, the user-defined probability distributions for the segment identifiers and the corresponding values represent exactly this underlying logic.
+Generating content with PIPCG effectively corresponds to drawing a sample from the random variable $\boldsymbol{C}\sim p(C)$, where
+$p(C)=p(v_{1},\dots,v_{N})$ (2)
+denotes the probability to generate the content $C$ in the form of Equation 1 with values $v_{1},\dots,v_{N}$.
+Hence, $\underline{\mathcal{C}}:=\{C\,|\,C\in\mathcal{C}\,\land\,p(C)>0\}$ represents the set of all content instances that can potentially be generated out of all possible content instances $\mathcal{C}$.
+Classical Wave Function Collapse (CWFC)
+CWFC can be understood as a special case of PIPCG.
+In this manuscript, only a simplified form of the simple tiled model of CWFC is considered (with the minor extension of graph-based adjacency [12]), as it is sufficient to capture the key components of the approach.
+To realize CWFC, the two randomized selection steps of PIPCG are chosen as follows:
+  1. 1.
+    For the segment identifier selection, the segments with the smallest Shannon entropy with respect to the available value options are chosen.
+  2. 2.
+    Value selection is determined by a pattern-based set of user-defined rules.
+Adjacency.
 For the pattern-based value selection, the segments are organized in such a way that they have a predefined adjacency relationship to each other that does not change.
 For example, consider that the pixels of an image constitute the segments.
-Choosing adjacency between the nearest neighbors leads to an adjacency configuration with four directions (right, up, left, down).
-Formally, each direction within a adjacency configuration is specified by an index $d\in[1,D]$, where $D$ denotes the total number of directions.
-The direction-based adjacency relationship between different segments is then defined by the coefficients $\alpha^{d}_{i,j}\in\mathbb{B}:=\{0,1\}$ for $i,j\in[1,N]$ and $d\in[1,D]$.
-If $\alpha^{d}_{i,j}=1$, segment $i$ is connected to segment $j$ in direction $d$; otherwise, it is not connected.
-For each direction $d$, these coefficients form an adjacency matrix $\alpha^{d}\in\mathcal{B}:=\mathbb{B}^{N\times N}$ of a directed graph with $N$ vertices that represent the segments.
-Hence, the segment adjacency is in fact an abstract concept that is not necessarily related to the visual form of the content.
-An example is shown in fig.
-1.
+Choosing adjacency between the nearest neighbors leads to an adjacency configuration with four directions (right, up, left, down) as shown in Figure 1.
 Figure 1: Exemplary nearest neighbor adjacency configuration for CWFC on a two-dimensional grid consisting of $N=9$ segments.
 The adjacency relationship for each of the four directions right ($d=1$), up ($d=2$), left ($d=3$), and down ($d=4$) can be represented as a directed graph with adjacency matrix $\alpha^{d}$.
-Based on the chosen adjacency configuration, a pattern-based value selection can be defined.
-For this purpose, a more general rule-based selection is introduced first and then the pattern-based selection is presented as a special case.
-For a rule-based value selection, consider a set of $m$ rules
-$R:=\{r_{1},\dots,r_{m}\}\in\mathcal{R},$ (12)
-from the set of of all possible rulesets $\mathcal{R}$, where each rule
-$r_{i}:=r_{i}(v,w):=(v,w)$ (13)
-for $i\in[1,m]$ consists of a value $v\in A$ and a weight function $w:=w(i,\alpha^{d},C^{\prime})\in\mathbb{R}_{\geq 0}$ with
-$w:[1,N]\times\mathcal{B}\times\overline{\mathcal{C}}\rightarrow\mathbb{R}_{\geq 0},$ (14)
-where
-$\overline{\mathcal{C}}:=\cup_{k=0}^{N-1}\mathcal{C}(k)$ (15)
-denotes the unified set of all partial contents. Given a ruleset $R\in\mathcal{R}$, the rule-based version of the probability distribution from eq. 6 is defined as
-$p_{k}^{\mathrm{v}}(\nu|i,C^{\prime})=p_{k}^{\mathrm{v,R}}(\nu|i,C^{\prime}):=\frac{F_{k}(\nu,i,C^{\prime})}{\overline{F}_{k}(i,C^{\prime})}$ (16)
-with the abbreviations
-$F_{k}(\nu,i,C^{\prime}):=\sum_{\begin{subarray}{c}(v,w)\in R\\ v=\nu\end{subarray}}w(i,\alpha^{d},C^{\prime})$ (17)
-and
-$\overline{F}_{k}(i,C^{\prime}):=\sum_{\nu\in A}F_{k}(\nu,i,C^{\prime}),$ (18)
-respectively, for $\nu\in A$, $i\in[1,N]$, $C^{\prime}\in\mathcal{C}(k)$ for $k\in[1,N]$, and $\alpha^{d}\in\mathcal{B}$ for $d\in[1,D]$.
-Summarized, the set of rules $R$ determines how values are selected based on the non-vanishing results of the respective weight functions.
-Note that condition V1 is always satisfied, whereas in case of $\overline{F}_{k}(i,C^{\prime})\overset{!}{=}0$ for any $C^{\prime}\in\mathcal{C}(k)$ with $p_{k}^{\mathrm{s,H}}(i|C^{\prime})p(C^{\prime})>0$, condition V2 is violated.
-Pattern-based value selection is a special case of rule-based value selection, where the weight function
-$w(i,\alpha^{d},C^{\prime})\!:=\!w^{\mathrm{P}}(i,\alpha^{d},C^{\prime},P,u)\!:=\!u\tau(i,\alpha^{d},C^{\prime},P)$ (19)
-of each rule
-$r_{i}:=r_{i}^{\mathrm{P}}(v,P,u):=(v,w^{\mathrm{P}}(i,\alpha^{d},C^{\prime},P,u))$ (20)
-for $i\in[1,m]$ is defined by a non-vanishing factor $u\in\mathbb{R}_{>0}$ and a pattern
-$P:=\{(d_{1},v_{1}),\dots,(d_{n},v_{n})\}\in\mathcal{P}$ (21)
-consisting of a set of $n\leq D$ direction-value pairs with $d_{i}\in[1,D]$ for $i\in[1,n]$, $d_{i}\neq d_{j}\,\forall\,i,j\in[1,n]\land i\neq j$, and $v_{i}\in A$ for $i\in[1,n]$, where
-$\mathcal{P}:=\{P\,|\,P\in([1,D]\times A)^{n}\,\forall\,n\in[0,D]\}$ (22)
-denotes the set of all possible patterns. A pattern $P$ is fulfilled, if
-$\tau(i,\alpha^{d},C^{\prime},P):=\prod_{(d,v)\in P}\prod_{\begin{subarray}{c}(s,v^{\prime})\in C^{\prime}\\ \alpha^{d}_{i,s}=1\end{subarray}}\delta_{vv^{\prime}}\in[0,1]$ (23)
-evaluates to one. Here, $\delta$ represents the Kronecker delta. The set of all possible pattern-based rulesets is denoted by $\mathcal{R}^{\mathrm{P}}\subset\mathcal{R}$.
-Summarized, the weight function, eq.
-19, yields $u$ if the pattern $P$ applies and zero otherwise.
-In line with expectations, only adjacent segments (i.e., $\alpha^{d}_{i,s}=1$) have an influence on eq.
-16 according to eq.
-23.
-Furthermore, the rules from eq. 20 can be rewritten as
+Formally, each direction within a adjacency configuration is specified by an index $d\in[1,D]$, where $D$ denotes the total number of directions.
+The direction-based adjacency relationship between different segments is then defined by the coefficients $\alpha^{d}_{i,j}\in\{0,1\}$ for $i,j\in[1,N]$ and $d\in[1,D]$.
+If $\alpha^{d}_{i,j}=1$, segment $i$ is connected to segment $j$ in direction $d$; otherwise not.
+For each direction $d$, these coefficients form an adjacency matrix $\alpha^{d}\in\{0,1\}^{N\times N}$ of a directed graph with $N$ vertices that represent the segments.
+Hence, the segment adjacency is in fact an abstract concept that is not necessarily related to the visual form of the content.
+Patterns.
+Patterns define how values are chosen based on the already generated content. A pattern
+$P:=\{(d_{1},v_{1}),\dots,(d_{n},v_{n})\}$ (3)
+consists of a set of $n\leq D$ direction-value pairs with $d_{i}\in[1,D]$ and $v_{i}\in A$ for $i\in[1,n]$, which represents a layout of segments. A pattern-based rule
 $r_{i}^{\mathrm{P}}(v,P,u)=(v,u,P)$ (4)
-for $i\in[1,m]$ with a value $v\in A$, a factor $u\in\mathbb{R}_{>0}$, and a pattern $P\in\mathcal{P}$. For the pattern-based value selection, eq. 16 can be reformulated as
-$p_{k}^{\mathrm{v,R}}(\nu|i,C^{\prime})$ $=p_{k}^{\mathrm{v,P}}(\nu|i,C^{\prime})$
-$:=\sum_{\begin{subarray}{c}(v,u,P)\in R\\ v=\nu\end{subarray}}\gamma(u,i,C^{\prime})\tau(i,\alpha^{d},C^{\prime},P)$ (27)
-with the coefficients
-$\gamma(u,i,C^{\prime}):=\frac{u}{\overline{F}_{k}(i,C^{\prime})}\in[0,1]$ (28)
-that correspond to the normalized weights of the rules.
-In the context of this manuscript, one variant (that lies between rule- and pattern-based selection) is particularly important, namely the use of a functional factor instead of a constant factor for pattern-based rules of the form of eq.
-24.
-Specifically, a factor
-$u:=u(i,C):[1,N]\times\overline{\mathcal{C}}\rightarrow\mathbb{R}$ (29)
-is considered that can depend both on the coordinate $i\in[1,N]$ and/or the already generated content $C\in\overline{\mathcal{C}}$, eq.
-15.
-By mapping to zero, this functional factor can disable a pattern for certain coordinates or partial contents.
-1.2.2 Identifier selection.
-For the entropic segment identifier selection, the probability distribution from eq. 5 is defined as
-$p_{k}^{\mathrm{s}}(i|C^{\prime})=p_{k}^{\mathrm{s,H}}(i|C^{\prime}):=\begin{cases}|I_{k}(C^{\prime})|^{-1}&\text{if }i\in I_{k}(C^{\prime})\\ 0&\text{otherwise}\end{cases}$ (30)
-with the set of optimal solutions
-$I_{k}(C^{\prime}):=\{i\,|\,i\in S^{-}(C^{\prime})\land H_{k}(i,C^{\prime})=h_{k}(C^{\prime})\}$ (31)
-of the optimization task
-$h_{k}(C^{\prime}):=\operatorname*{arg\,min}_{j\in S^{-}(C^{\prime})}H_{k}(j,C^{\prime}),$ (32)
-where
-$H_{k}(i,C^{\prime}):=-\sum_{v\in A}$ $p_{k}^{\mathrm{v,P}}(\nu=v|i,C^{\prime})$
-×$\times\ln p_{k}^{\mathrm{v,P}}(\nu=v|i,C^{\prime})$ (33)
-denotes the Shannon information entropy for $C^{\prime}\in\mathcal{C}(k-1)$, $k\in[1,N]$ That is, the segments with the smallest Shannon entropy with respect to the available value options are chosen.
-1.3 CWFC Example: Checkerboard
+is defined by a value $v\in A$, a factor $u\in\mathbb{R}_{>0}$, and a pattern $P$.
+It describes the weight $u$ of selecting a value $v$ for the target segment given the layout of adjacent segments $P$.
+The higher the value $u$, the more probable becomes the selection of the corresponding value.
+A selection mechanism with $m$ rules is then fully defined by a ruleset
+$R:=\{r_{1},\dots,r_{m}\}$ (5)
+comprised of rules from Equation 4 with $i\in[1,m]$.
+CWFC Example: Checkerboard
 As a simple example, consider the generation of images with black and white checkerboard patterns.
 The pixels of an image constitute the segments and the alphabet consists of only $W=2$ symbols standing for the two colors.
-As an adjacency configuration, the nearest-neighbor adjacency from fig.
-1 is presumed.
-A checkerboard image can be achieved with only two patterns of the form of eq.
-21, namely $P_{1}:=\{(1,1),(2,1),(3,1),(4,1)\}$ and $P_{2}:=\{(1,2),(2,2),(3,2),(4,2)\}$.
-The corresponding ruleset, eq.
-12, reads $R=\{r_{1}^{\mathrm{P}},r_{2}^{\mathrm{P}}\}$ and therefore contains the $m=2$ rules $r_{1}^{\mathrm{P}}(v=2,P=P_{1},u=1)$ and $r_{2}^{\mathrm{P}}(v=1,P=P_{1},u=1)$ based on eq.
-20, see fig.
-2.
+As an adjacency configuration, the nearest-neighbor adjacency from Figure 1 is presumed.
+A checkerboard image can be achieved with only two patterns of the form of Equation 3, namely $P_{1}:=\{(1,1),(2,1),(3,1),(4,1)\}$ and $P_{2}:=\{(1,2),(2,2),(3,2),(4,2)\}$.
+The corresponding ruleset, Equation 5, reads $R=\{r_{1}^{\mathrm{P}},r_{2}^{\mathrm{P}}\}$ and therefore contains the $m=2$ rules $r_{1}^{\mathrm{P}}(v=2,P=P_{1},u=1)$ and $r_{2}^{\mathrm{P}}(v=1,P=P_{1},u=1)$, see Figure 2.
 Figure 2: Visualization of pattern-based rules.
 (a) Visualization scheme.
-The center tile corresponds to the target segment $i$ of the weight function, eq.
-19, with a target value of $v$, whereas the adjacent tiles represent the required values $v_{1},\dots,v_{4}\in\{1,2\}$ of the four adjacent segments in the respective directions $d=1,\dots,d=4$.
-(b) Pattern-based rules $r_{1}^{\mathrm{P}}(v=2,P=P_{1},u=1)$ and $r_{2}^{\mathrm{P}}(v=1,P=P_{2},u=1)$, eq.
-20, that can be used for the generation of images with black and white checkerboard patterns.
-The iterative process of CWFC is sketched in fig.
-3 for a $3\times 3$ image (represented by $N=9$ segments).
-Only two content instances can be generated, $C_{1}:=\{(1,1),\dots\}$ and $C_{2}:=\{(1,2),\dots\}$ with equal probability, eq.
-10.
+The center tile corresponds to the target segment $i$ with a target value of $v$, whereas the adjacent tiles represent the required values $v_{1},\dots,v_{4}\in\{1,2\}$ of the four adjacent segments in the respective directions $d=1,\dots,d=4$.
+(b) Pattern-based rules $r_{1}^{\mathrm{P}}(v=2,P=P_{1},u=1)$ and $r_{2}^{\mathrm{P}}(v=1,P=P_{2},u=1)$, Equation 4, that can be used for the generation of images with black and white checkerboard patterns.
+The iterative process of CWFC is sketched in Figure 3 for a $3\times 3$ image (represented by $N=9$ segments).
+Only two content instances can be generated, $C_{1}:=\{(1,1),\dots\}$ and $C_{2}:=\{(1,2),\dots\}$ with equal probability $p(C_{1})=p(C_{2})=\frac{1}{2}$, Equation 2.
 The choice of the first value $v_{s(1)}$ already determines the choices of all other values.
-Figure 3: CWFC for the generation of $3\times 3$ images with checkerboard patterns defined by the rules from fig.
-2.
-In each iteration $k\in[1,9]$, three steps take place: (i) a segment identifier $s(k)$ is drawn, eq.
-30, (ii) a corresponding value $v_{s(k)}$ is drawn, section 1.2.1, and (iii) the newly generated identifier-value pair is added as a segment to the content instance $C(k)$.
+Figure 3: CWFC for the generation of $3\times 3$ images with checkerboard patterns defined by the rules from Figure 2.
+In each iteration $k\in[1,9]$, three steps take place: (i) a segment identifier $s(k)$ with the smallest entropy is drawn, (ii) a corresponding value $v_{s(k)}$ is drawn according to the pattern-based ruleset, and (iii) the newly generated identifier-value pair is added as a segment to the content instance $C(k)$.
 For $s(k)$ and $v(k)$, the uniformly distributed set of possible choices are listed.
-The choices for $s(k)$ depend on the Shannon entropy $H_{k}$, section 1.2.2, of each undefined segment, as shown on the right.
+The choices for $s(k)$ depend on the Shannon entropy $H_{k}$ of each undefined segment (see appendix), as shown on the right.
 The choices for $v(k)$ depend on the fulfillment of the two rules $r_{1}^{\mathrm{P}}$ and $r_{2}^{\mathrm{P}}$, as listed.
-2 QUANTUM METHODS
-Based on the previously established classical concepts, a QWFC method and a HWFC method are proposed in the following. A software implementation is available online [9].
-2.1 Quantum Wave Function Collapse (QWFC)
-So far, the procedural generation concepts have been presented from a purely classical perspective.
-However, it seems natural that the probability distribution of the content instances $p(C)$, eq.
-10, can also be realized with the help of a quantum circuit by utilizing the intrinsic randomness of quantum physics [10].
-This is the very idea on which QWFC is based.
-A single alphabet $A$ with $W$ symbols is presumed, eq. 11. Furthermore, two procedures control the content generation, namely segment identifier selection and value selection.
-2.1.1 Identifier selection.
-The Identifier selection is performed with a predefined order given by the vector $\sigma\in S_{N}$, which is the set of all permutations of $[1,N]$. Hence,
-$p_{k}^{\mathrm{s}}(i|C^{\prime})=p_{k}^{\mathrm{s,o}}(i|C^{\prime}):=\delta_{i\sigma_{i}}.$ (34)
-This approach leads to (partial) content instances $\mathcal{C}(k)=\mathcal{C}_{\sigma}(k)$, eq. 2, of the form
-$\mathcal{C}_{\sigma}(k):=\{C\,|\,$ $C=\{(\sigma_{1},v_{\sigma_{1}}),\dots,(\sigma_{k},v_{\sigma_{k}})\}$
-$\,\forall\,v_{\sigma_{l}}\in A\,\forall\,l\in[1,k]\}$ (35)
-for $k\in[1,N]$ and $\mathcal{C}_{\sigma}(0):=\{\{\}\}$.
-2.1.2 Value selection.
-The value selection is pattern-based in analogy to section 1.2.1. In particular, the value for each segment is encoded by a set of qubits, which requires
-$q:=\lceil\log_{2}W\rceil$ (36)
-qubits for each segment $i\in[1,N]$. The group of $q$ qubits that represent the value for segment $i$ is denoted by
-$\mathcal{Q}_{i}:=\{q_{i}^{1},\dots,q_{i}^{q}\},$ (37)
-where $q_{i}^{j}$ references the $j$th qubit of this group for $j\in[1,q]$.
-The state of these qubits encodes the value $v_{i}$ of the segment $i$ in binary as the tensor product sequence of qubit states
-$\bigotimes_{j=1}^{q}\ket{b_{j}}=\Ket{v_{i}-1=\sum_{j=1}^{q}b_{j}2^{j-1}}_{i}=\ket{v_{i}-1}_{i}\in\mathcal{H}_{i}$ (38)
-with $b_{j}\in\mathbb{B}$, $v_{i}\in A$, and $i\in[1,N]$, where $\mathcal{H}_{i}$ denotes the joint Hilbert space of all qubit from $\mathcal{Q}_{i}$.
-In total,
-$Q:=Nq$ (39)
-qubits are required to represent the value of all $N$ segments.
-The joint Hilbert space of all qubits from $\bigcup_{i=1}^{N}\mathcal{Q}_{i}$ is given by $\mathcal{H}:=\bigotimes_{i=1}^{N}\mathcal{H}_{i}$ and a state |$\ket{\Psi}\in\mathcal{H}$ can realize samples from a specific probability distribution of content instances
-$p(C\!=\!\{(1,v_{1}),\dots,(N,v_{N})\})=|\bra{\Psi}\bigotimes_{i=1}^{N}\ket{v_{i}-1}_{i}|^{2}.$ (40)
-based on a projective measurement.
-After these preliminary considerations, the remaining question is how such a state can be prepared constructively with quantum gates from a given pattern-based ruleset.
-This question will be answered in the following.
-2.1.3 Quantum circuit.
-Initially, all qubits are prepared in the ground state, i.e., |$\ket{\Psi}=\ket{0}:=\bigotimes_{i=1}^{N}\ket{0}_{i}$.
-Then, given a pattern-based ruleset $R\in\mathcal{R}^{\mathrm{P}}$, eq.
-12, a series of operators is applied, one for each iteration $k\in[1,N]$ such that |$\ket{\Psi}:=U_{N}\cdots U_{1}\ket{0}$ represents the final state after $k$ iterations.
-In each iteration $k$, the state of the qubits $\mathcal{Q}_{\sigma_{k}}$ (that represent the value for segment $\sigma_{k}$) is prepared conditioned on the state of (a subset of) the qubits $\mathcal{Q}_{\sigma_{1}},\dots,\mathcal{Q}_{\sigma_{k-1}}$ (that represent the values of the already prepared segments $\sigma_{1},\dots,\sigma_{k-1}$), which leads to an entangled joint state.
-A sketch is shown in fig.
-4 (a).
+QUANTUM METHODS
+In the previous section, only classical PCG was considered.
+This section is about QPCG.
+First, QWFC is proposed as a special case of PIPCG.
+For this purpose, the probability distribution of the content instances $p(C)$, Equation 2, is realized with the help of a quantum circuit and the sampling is performed by exploiting the intrinsic randomness of quantum measurements.
+That is, the actual quantum mechanical wave function collapse is used to implement a quantum version of the Wave Function Collapse algorithm.
+Subsequently, this section proposes HWFC as a quantum-classical hybrid algorithm based on content partitioning to overcome the limitations of NISQ devices.
+Only the main features of QWFC and HWFC are summarized, the technical details can be found in the appendix.
+A software implementation is available online [22].
+Quantum Wave Function Collapse (QWFC)
+To realize QWFC, the two randomized selection steps of PIPCG are chosen as follows:
+  1. 1.
+    The segment identifier selection is performed with a predefined order given by the vector $\sigma\in S_{N}$, where $S_{N}$ is the set of all permutations of $[1,N]$.
+  2. 2.
+    The value selection is based on a pattern-based ruleset in the same way as for CWFC.
+Value Encoding.
+The value for each segment is encoded by a set of qubits, which requires $q:=\lceil\log_{2}W\rceil$ qubits for each segment $i\in[1,N]$.
+The group of $q$ qubits that represent the value for segment $i$ is denoted by $\mathcal{Q}_{i}:=\{q_{i}^{1},\dots,q_{i}^{q}\}$, where $q_{i}^{j}$ references the $j$th qubit of this group for $j\in[1,q]$.
+The state of these qubits encodes the value $v_{i}$ of the segment $i$ in binary as the tensor product sequence of qubit states $\bigotimes_{j=1}^{q}\ket{b_{j}}=\ket{v_{i}-1}_{i}\in\mathcal{H}_{i}$ via $v_{i}-1=\sum_{j=1}^{q}b_{j}2^{j-1}$ with $b_{j}\in\{0,1\}$.
+Here, $\mathcal{H}_{i}$ denotes the joint Hilbert space of all qubit from $\mathcal{Q}_{i}$ with $v_{i}\in A$ for $i\in[1,N]$.
+In total, $Q:=Nq$ qubits are required to represent the value of all $N$ segments.
+The joint Hilbert space of all qubits from $\bigcup_{i=1}^{N}\mathcal{Q}_{i}$ is given by $\mathcal{H}:=\bigotimes_{i=1}^{N}\mathcal{H}_{i}$.
+Probability Encoding.
+Presume a state |$\ket{\Psi}\in\mathcal{H}$ that can realize samples from $p(C)$, Equation 2, with
+$p(C\!=\!\{(1,v_{1}),\dots,(N,v_{N})\})=\big{|}\bra{\Psi}\bigotimes_{i=1}^{N}\ket{v_{i}-1}_{i}\big{|}^{2}.$ (6)
+To prepare |$\ket{\Psi}$ from a ground state |$\ket{0}:=\bigotimes_{i=1}^{N}\ket{0}_{i}$ based on a pattern-based ruleset $R$, Equation 5, a series of operators is applied, one for each iteration $k\in[1,N]$ such that |$\ket{\Psi}:=U_{N}\cdots U_{1}\ket{0}$ represents the final state after $k$ iterations.
+In each iteration $k$, the state of the qubits $\mathcal{Q}_{\sigma_{k}}$ (that represent the value for segment $\sigma_{k}$) is prepared conditioned on the state of (a subset of) the qubits $\mathcal{Q}_{\sigma_{1}},\dots,\mathcal{Q}_{\sigma_{k-1}}$ (that represent the values of the already prepared segments $\sigma_{1},\dots,\sigma_{k-1}$), which leads to an entangled joint state as sketched in Figure 4(a).
 Figure 4: QWFC circuit with $\sigma=(1,\dots,N)$.
-(a) Circuit layout with an exemplarily alphabet $A$, eq.
-11, with $W=16$ symbols, which requires a set $\mathcal{Q}_{i}$, eq.
-37, of $q=4$ qubits, eq.
-36, for each segment with $i\in[1,N]$.
-In each iteration $k\in[1,N]$, the operator $U_{k}$, eq.
-41, prepares the qubits from $\mathcal{Q}_{k}$ in a superposition state |$\ket{C^{\prime}}_{k}$, eq.
-45, entangled with (a subset of) the qubits from $\mathcal{Q}_{1},\dots,\mathcal{Q}_{k-1}$.
-All qubits are measured to obtain $p(C)$, eq.
-40.
-(b) Gate decomposition of $U_{k}$ into the components of $U_{k}^{\mathrm{coin}}$, eq.
-42, which consists of pairs $U_{k}^{\mathrm{ctrl}}(C^{\prime})\otimes U_{k}^{\mathrm{init}}(C^{\prime})$, eqs.
-43 and 44.
-Each of these pairs represent a conditional loading of a probability distribution to realize |$\ket{C^{\prime}}_{k}$ (which is not explicitly shown).
-The combined control operators act on the qubits from $\mathcal{Q}_{k}^{+}$ eq.
-47, whereas the initialization operators act on the qubits from $\mathcal{Q}_{k}$.
-Specifically,
-$U_{k}:=\left[U_{k}^{\mathrm{coin}}+U_{k}^{0}\right]\otimes\bigotimes_{l=k+1}^{N}\mathbbm{1}_{\sigma_{l}}\in\mathcal{H}$ (41)
-for $k\in[1,N]$ contains $\mathbbm{1}_{i}\in\mathcal{H}_{i}$, the unit operator for $i\in[1,N]$ and two other operators. First, the conditional initialization operator
-$U_{k}^{\mathrm{coin}}=\sum_{C^{\prime}\in\mathcal{C}_{\sigma}(k)}\left[U_{k}^{\mathrm{ctrl}}(C^{\prime})\otimes U_{k}^{\mathrm{init}}(C^{\prime})\right],$ (42)
-where $\mathcal{C}_{\sigma}(k)$ represents the set of possible partial content instances from section 2.1.1,
-$U_{k}^{\mathrm{ctrl}}(C^{\prime}):=\bigotimes_{(s,v)\in L_{k}(C^{\prime})}\ket{v-1}\bra{v-1}_{s}\in\bigotimes_{l=1}^{k-1}\mathcal{H}_{\sigma_{l}}$ (43)
-denotes the control operator, and
-$U_{k}^{\mathrm{init}}(C^{\prime}):=\ket{C^{\prime}}\bra{0}_{\sigma_{k}}\in\mathcal{H}_{\sigma_{k}}$ (44)
-the initialization operator with
-|$\ket{C^{\prime}}_{\sigma_{k}}:=\sum_{\nu\in A}\sqrt{p_{k}^{\mathrm{v,P}}(\nu|\sigma_{k},C^{\prime})}\ket{\nu-1}_{\sigma_{k}}\in\mathcal{H}_{\sigma_{k}},$ (45)
-which is based on the pattern-based probability distribution of values $p_{k}^{\mathrm{v,P}}(\nu|i,C^{\prime})$ from section 1.2.1. The summation in eq. 43 runs over all adjacent segments
-$L_{k}(C^{\prime}):=\{(s,v)\,|\,$ $(s,v)\in C^{\prime}\,\land\,\exists\,P\in R_{P}$
-$\exists\,d\in P_{d}:\alpha^{d}_{\sigma_{k}s}=1\}\subseteq C^{\prime},$ (46)
-where $R_{P}:=\{P_{1},\dots,P_{m}\}$ stands for the set of patterns in the pattern-based ruleset $R\in\mathcal{R}^{\mathrm{P}}$ with $P_{i}\in\mathcal{P}$ for $i\in[1,m]$ and $P_{d}:=\{d_{1},\dots,d_{n}\}$ stands for the set of directions in the pattern $P\in\mathcal{P}$, eq.
-21, with $d_{j}\in[1,D]$ for $j\in[1,n]$.
-In particular, as discussed above for the classic case, choosing $C^{\prime}$ instead of $L_{k}(C^{\prime})$ in the summation has no effect on $p_{k}^{\mathrm{v,P}}(\nu|i,C^{\prime})$ and thus |$\ket{C^{\prime}}_{\sigma_{k}}$.
-In iteration $k$, only the qubits in $\mathcal{Q}_{k-1}^{+}$ with
-$\mathcal{Q}_{k}^{+}:=\{q\,|\,q=q_{k}^{i}\in\mathcal{Q}_{\sigma_{k}}\,\forall\,i\in[1,q]\,\land\,\sigma_{k}\in L_{k}^{+}\}$ (47)
-are affected by $U^{\mathrm{ctrl}}(C^{\prime})$ for some $C^{\prime}\in\mathcal{C}_{\sigma}(k)$, where
-$L_{k}^{+}:=\{s\,|\,\exists\,C^{\prime}\in\mathcal{C}_{\sigma}(k)\,:\,s\in S^{+}(L_{k}(C^{\prime}))\}$ (48)
-denotes the affected segments for $k=[1,N]$. Here, $\mathcal{Q}_{k}^{+}\subseteq\{\mathcal{Q}_{1},\dots,\mathcal{Q}_{k}\}$ and $L_{k}^{+}\subseteq\{\sigma_{1},\dots,\sigma_{k}\}$, respectively.
-The second operator in eq. 41 is the neutral operator
-$U_{k}^{0}=\sum_{C^{\prime}\in\mathcal{C}_{\sigma}(k)}\left[U_{k}^{\mathrm{ctrl}}(C^{\prime})\otimes U_{k}^{\mathrm{init},0}+U_{k}^{\mathrm{ctrl},0}(C^{\prime})\right].$ (49)
-Initially, all qubits $\mathcal{Q}_{k}$ are in the ground state. Therefore,
-$U_{k}^{\mathrm{init},0}:=\sum_{v\in[2,W]}\ket{v-1}\bra{v-1}_{\sigma_{k}}\in\mathcal{H}_{\sigma_{k}}$ (50)
-vanishes. Furthermore,
-$U_{k}^{\mathrm{ctrl},0}(C^{\prime})\!:=\!\left[\bigotimes_{l=1}^{k-1}\mathbbm{1}_{\sigma_{l}}-U_{k}^{\mathrm{ctrl}}(C^{\prime})\right]\otimes\mathbbm{1}_{\sigma_{k}}\!\in\bigotimes_{l=1}^{k}\mathcal{H}_{\sigma_{l}}$ (51)
-does not change the state by construction.
-The operator $U_{k}$, eq.
-41, can be straightforwardly realized with quantum gates for a given $k\in[1,N]$.
-To this end, only the conditional initialization operator $U_{k}^{\mathrm{coin}}$, eq.
-42, has to be considered because all other parts of $U_{k}$ have no effect on the resulting state.
-The conditional initialization operator consist of a sequence of joint pairs of control-initialization operators $U_{k}^{\mathrm{ctrl}}(C^{\prime})\otimes U_{k}^{\mathrm{init}}(C^{\prime})$, one for each $C^{\prime}\in\mathcal{C}_{\sigma}(k)$.
-Each of these joint pairs represent a conditional loading of a probability distribution to realize |$\ket{C^{\prime}}_{\sigma_{k}}$, which is a well-known task that may, however, require an exponential number of gates [11, 12, 13].
-A sketch is shown in fig.
-4 (b).
-For the definition of QWFC, a noise-free, ideal quantum device is presumed, that is able to run the emerging circuits without any errors.
-However, the currently available noisy intermediate-scale quantum (NISQ) hardware exhibits significant noise and uncertainty [14].
-As a consequence, the content sampled from eq.
-40 can also be expected to violate the prescribed patterns.
-This will be demonstrated further below for several PCG use cases.
-The proposed conditions S1, S2, V1, and V2 are in fact not mandatory in the original implementation of CWFC, where a violation leads to a restart of the algorithm.
-This can also be considered in QWFC by adding an additional “conflict detection qubit” that stores this information, an extension that is not further discussed here.
-2.2 QWFC Example: Checkerboard
-In the following, the above example for the generation of $3\times 3$ checkerboard images is considered for QWFC, where the predefined segment order $\sigma=(1,2,3,6,5,4,7,8,9)$ is used for eq.
-34 as visualized in fig.
-5 (a).
-The resulting CWFC circuit is shown in fig.
-5 (b) and the resulting probability distribution of content instances $p(C)$, eq.
-40, in fig.
-5 (c).
-Given a noise-free quantum device, a measurement of the circuit corresponds to drawing a sample from this distribution.
-Figure 5: QWFC for the generation of $3\times 3$ images with black and white checkerboard patterns in analogy to fig.
-3.
-(a) Predefined segment order $\sigma=(1,2,3,6,5,4,7,8,9)$ for eq.
-34.
-(b) Circuit layout with the same symbols as in fig.
-4.
-(c) Generated content instances $C_{1}$ and $C_{2}$ with $p(C_{1})=p(C_{2})=\frac{1}{2}$, eq.
-40.
-2.3 Hybrid Quantum-Classical WFC (HWFC)
-The required number of qubits for the proposed QWFC method increases linearly with the number of segments and logarithmically with the number of symbols in the alphabet as shown in eqs.
-36 and 39.
-At the same time, the respective circuits may become exponentially deeper because of the costly conditional preparation of probability distributions.
-The reliable execution of such big circuits on NISQ hardware can be very challenging or even impossible.
+(a) Circuit layout with an exemplary alphabet $A$ with $W=16$ symbols, which requires a set $\mathcal{Q}_{i}$ of $q=4$ qubits for each segment with $i\in[1,N]$.
+In each iteration $k\in[1,N]$, the operator $U_{k}$ prepares the qubits from $\mathcal{Q}_{k}$ in a superposition state |$\ket{C^{\prime}}_{k}$, entangled with (a subset of) the qubits from $\mathcal{Q}_{1},\dots,\mathcal{Q}_{k-1}$.
+All qubits are measured to obtain $p(C)$ (see appendix).
+(b) Gate decomposition of $U_{k}$ into the components of $U_{k}^{\mathrm{coin}}$, which consists of pairs $U_{k}^{\mathrm{ctrl}}(C^{\prime})\otimes U_{k}^{\mathrm{init}}(C^{\prime})$.
+Each of these pairs represent a conditional loading of a probability distribution.
+The combined control operators act on the qubits from $\mathcal{Q}_{k}^{+}$, whereas the initialization operators act on the qubits from $\mathcal{Q}_{k}$ (see appendix).
+The required number of qubits for the proposed QWFC method increases linearly with the number of segments and logarithmically with the number of symbols in the alphabet.
+At the same time, the respective circuits may become exponentially deeper because of the costly conditional preparation of probability distributions [23].
+QWFC Example: Checkerboard
+In the following, the above example for the generation of $3\times 3$ checkerboard images is considered for QWFC, where the predefined segment order $\sigma=(1,2,3,6,5,4,7,8,9)$ is used as visualized in Figure 5(a).
+The resulting CWFC circuit is shown in Figure 5(b) and the resulting probability distribution of content instances $p(C)$, Equation 6, in Figure 5(c).
+A (noise-free) measurement of the circuit corresponds to drawing a sample from this distribution.
+Figure 5: QWFC for the generation of $3\times 3$ images with black and white checkerboard patterns in analogy to Figure 3.
+(a) Predefined segment order $\sigma=(1,2,3,6,5,4,7,8,9)$.
+(b) Circuit layout with the same symbols as in Figure 4.
+(c) Generated content instances $C_{1}$ and $C_{2}$ with $p(C_{1})=p(C_{2})=\frac{1}{2}$.
+Hybrid Quantum-Classical WFC (HWFC)
 […]
-Finally, the distribution of the joint content instances is given by
-$p\left(C=\bigcup_{h=1}^{H}C^{h}\right)=\prod_{h=1}^{H}p^{h}(C^{h}|C^{1},\dots,C^{h-1}),$ (52)
-which corresponds to performing a QWFC on the complete content, cf.
-fig.
-6.
-With the proposed HWFC method, the size of the QWFC circuits can be reduced at the price of an additional outer iteration loop and the inclusion of classical correlations in the conditional QWFC circuit generation.
 Figure 6: HWFC with $H=3$ partitions of $N=9$ segments. For each partition, a conditional QWFC is performed and the resulting content instances are combined.
-3 DEMONSTRATION
-In the present section, five simple content creation use cases are presented to demonstrate the proposed QWFC and HWFC methods.
-For the realization of these examples, both quantum computing simulators and actual IBM Quantum devices were used, which were accessed via the IBM Quantum cloud services [15] during December 2023.
-3.1 Checkerboard Revisited
-As a first use case, the previously presented checkerboard example is recalled, cf.
-figs.
-2 and 5.
+DEMONSTRATION
+In the present section, five simple content creation use cases are presented to demonstrate the proposed methods.
+For QWFC and HWFC, both idealized (i.e., noise-free) simulators and IBM Quantum devices were used, the latter being accessed via the IBM Quantum Cloud Services [24] during December 2023.
+Checkerboard Revisited
+The first use case is the previously presented checkerboard example, cf.
+Figures 2 and 5.
 Again, $3\times 3$ instances are considered, for which a comparison between QWFC and HWFC is performed on the IBM Quantum device ibm_kyoto (Eagle r3 processor version 1.2.4, 127 qubits).
 For QWFC, 10 0001000010\,00010 000 shots are measured, each yielding a content instance.
-For HWFC, $H=3$ equal-sized partitions are used in analogy to fig.
-6 and the method is repeated $131$ times, each time yielding a content instance.
-The results are shown in fig.
-7.
+For HWFC, $H=3$ equal-sized partitions are used in analogy to Figure 6 and the method is repeated $131$ times, each time yielding a content instance.
+The results are shown in Figure 7.
 […]
-In total, there are $5122^{9}=$ 512 possible content instances $C\in\mathcal{C}$, eq.
-1, but only two of them, $C_{1}$ and $C_{2}$, are valid.
-The remaining $510$ content instances are invalid because they violate the patterns.
-In fig.
-7 (a), QWFC is able to generate the two valid instances, but at the same time also produces a lot of invalid instances due to hardware imperfections.
-From fig.
-7 (b) it can be seen that the results from HWFC are much closer to the idealized solution that would contain only $C_{1}$ and $C_{2}$ with equal probability.
+In total, there are $5122^{9}=$ 512 possible content instances $C\in\mathcal{C}$, but only two of them, $C_{1}$ and $C_{2}$, are valid.
+The remaining $510$ content instances are invalid because they violate the prescribed patterns.
+In Figure 7(a), QWFC is able to generate the two valid instances, but at the same time also produces a lot of invalid instances due to hardware imperfections.
+From Figure 7(b) it can be seen that the results from HWFC are much closer to the idealized solution that would contain only $C_{1}$ and $C_{2}$ with equal probability.
 It is a consequence of the reduced circuit size from this method that invalid solutions occur with much less frequency than for QWFC.
-3.2 Pipes
+Pipes
 The second use case is to create an image from tiles of seven different pipe (or line) sections as well as one blank tile.
 That is, the alphabet consists of $W=8$ symbols and each tile in the image constitutes a segment.
-The alphabet is shown in fig.
-8 (a).
-As adjacency configuration, the nearest-neighbor adjacency from fig.
-1 is presumed with $D=4$.
+The alphabet is shown in Figure 8(a).
+As adjacency configuration, the nearest-neighbor adjacency from Figure 1 is presumed with $D=4$.
 The pattern-based ruleset is chosen in such a way that neighboring tiles have to form a connecting network of pipes.
-Three example rules for this purpose are shown in fig.
-8 (b).
+Three example rules for this purpose are shown in Figure 8(b).
 In total, $m=2048$ rules are required to achieve this goal.
@@
 (b) Example of three pattern-based rules $r_{1}^{\mathrm{p}}$, $r_{2}^{\mathrm{p}}$, and $r_{3}^{\mathrm{p}}$, cf.
-fig.
-2.
-(c) Example $10\times 4$ tile images that were generated using HWFC with $H=10$ on the IBM Quantum device ibm_kolkata and on simulator as well as CWFC.
+Figure 2.
+(c) Example $10\times 4$ tile images that were generated using HWFC with $H=10$ on the IBM Quantum device ibm_kolkata and a simulator; also shown are images from CWFC.
 Exemplarily, images consisting of $10\times 4$ tiles ($N=40$) are considered, for which a comparison between HWFC on the IBM Quantum device ibm_kolkata (Falcon r5.11 processor version 1.14.8, 27 qubits), HWFC on a simulator and CWFC is performed.
 For HWFC, $H=10$ equal-sized partitions are used, each corresponding to one column of the image.
-A few resulting instances are shown in fig.
-8 (c).
+A few resulting instances are shown in Figure 8(c).
 Based on a visual comparison, both CWFC and HWFC on the simulator yield similar results, as expected.
@@
 This demonstrates the limitations of the proposed method on NISQ devices.
-3.3 Hexagon Map
+Hexagon Map
 The third use case addresses images based on hexagonal tiles that can be interpreted as a map consisting of different terrains.
@@
 As adjacency configuration, a nearest-neighbor adjacency is presumed ($D=6$).
-The prescribed pattern-based ruleset only allows connections of the form blue-yellow-green-gray, which can be fulfilled with $m=1586$ rules, see fig.
-9 (a).
-To promote a higher occurrence of blue tiles, the factors in eq.
-24 are chosen as $u=5$ for blue tiles and $u=1$ otherwise.
+The prescribed pattern-based ruleset only allows connections of the form blue-yellow-green-gray, which can be fulfilled with $m=1586$ rules, see Figure 9(a).
+To promote a higher occurrence of blue tiles, the factors in Equation 4 are chosen as $u=5$ for blue tiles and $u=1$ otherwise.
 Figure 9: Map generation.
@@
 As a first example, images consisting of $N=7$ tiles are considered, for which QWFC on a simulator and CWFC are performed.
-The resulting distributions of content instances $p(C)$ in fig.
-9 (b) is exact for QWFC, eq.
-40, and based on 10 0001000010\,00010 000 samples for CWFC, eq.
-10.
-Without surprise, both distributions are significantly different.
+The resulting distributions of content instances $p(C)$ in Figure 9(b) is exactly calculated for QWFC, Equation 6, and based on 10 0001000010\,00010 000 samples for CWFC, Equation 2.
+Both distributions are significantly different, which is no surprise since the two approaches use a different segment identifier selection.
 The second example considers images from $N=127$ tiles, for which HWFC with $H=20$ is performed on a simulator.
-Example instances are shown in fig.
-9 (c).
-3.4 Platformer
+Example instances are shown in Figure 9(c).
+Platformer
 The fourth use case is motivated by computer game level design.
 The goal is to create a platformer-type level from a set of eight tiles [25], each corresponding to a different game element (ground, grass, mushroom, block, air, and a tree that is composed of three tiles).
-Consequently, the alphabet consists of $W=8$ symbols, see fig.
-10 (a).
+Consequently, the alphabet consists of $W=8$ symbols, see Figure 10(a).
 As adjacency configuration, the two tiles above and below are taken into account as neighbors ($D=2$).
-A pattern-based ruleset is prescribed that meets five requirements: (i) The bottom row consists of ground tiles and ground tiles can only be placed on top of each other, (ii) only a grass tile or a mushroom tile can be placed above a ground tile, (iii) the tree tiles must be placed in order with the lowest tile above a grass tile, (iv) air tiles can only be placed above grass, mushroom or other air tiles, (v) block tiles can only be placed between two air tiles.
-These requirements can be fulfilled with $m=15$ rules, see fig.
-10 (b).
-The first requirement can be resolved with a functional factor, eq.
-29, that vanishes for all non-ground tiles in the bottom row.
-To reduce the occurrence of block tiles, the corresponding factor in eq.
-24 is chosen as $u=$0.1$$ for $r_{6}^{\mathrm{p}}$.
-All other factors are set to $u=1$.
+A pattern-based ruleset is prescribed that meets five requirements: (i) The bottom row consists of ground tiles and ground tiles can only be placed on top of each other, (ii) only a grass tile or a mushroom tile can be placed above a ground tile, (iii) tree tiles must be placed in order with the lowest tile above a grass tile, (iv) air tiles can only be placed above grass, mushroom or other air tiles, (v) block tiles can only be placed between two air tiles.
+These requirements can be fulfilled with $m=15$ rules, see Figure 10(b).
+The first requirement can be resolved with a functional factor $u(i)$ in Equation 4 that vanishes for all non-ground tiles in the bottom row (see appendix).
+For all non-vanishing cases, $u=$0.1$$ for $r_{6}^{\mathrm{p}}$ and $u=1$ otherwise to reduce the occurrence of block tiles.
 The segment order $\sigma$ is chosen such that the levels are built up from bottom to top.
 Figure 10: Level generation.
-(a) Alphabet consisting of eight symbols, each of which represents a tile [16].
+(a) Alphabet consisting of $W=8$ symbols, each of which represents a tile [25].
 (b) Pattern-based ruleset consisting of $m=15$ rules.
@@
 Exemplarily, levels with $10\times 10$ tiles are considered, for which HWFC with $H=20$ equal-sized partitions (each corresponding to a half row of the level) is performed on a simulator.
-Four resulting instances are shown in fig.
-10 (c).
-3.5 Voxel Skyline
-The final use case considers the creation of a 3d voxel graphic with a binary alphabet of $W=2$ symbols, which represent the presence or absence of a voxel.
+Four resulting instances are shown in Figure 10(c).
+Voxel Skyline
+The final use case considers the creation of a three-dimensional voxel graphic with a binary alphabet of $W=2$ symbols, which represent the presence or absence of a voxel.
 Each voxel constitutes a segment, as adjacency configuration the voxel above and below are taken into account ($D=2$).
-The chosen pattern-based ruleset with $m=4$ ensures that voxels are built from the ground up, see fig.
-11 (a).
+The chosen pattern-based ruleset with $m=4$ ensures that voxels are built from the ground up, see Figure 11(a).
 This leads to a skyline-like structure.
-Figure 11: Skyline use case.
+Figure 11: Skyline generation.
 (a) Pattern-based ruleset consisting of $m=4$ rules in an analogous representation as in the previous figures.
@@
 (c) Example images from a $10\times 10\times 5$ voxel grid that were generated using QWFC on a simulator.
-Exemplarily, two cases are considered: First, images from a $4\times 4\times 4$ voxel grid, for which QWFC is performed both on a the IBM Quantum device ibm_osaka (Eagle r3 processor version 1.0.3, 127 qubits) and a simulator, see fig.
-11 (b).
+Exemplarily, two cases are considered: First, images from a $4\times 4\times 4$ voxel grid, for which QWFC is performed both on a the IBM Quantum device ibm_osaka (Eagle r3 processor version 1.0.3, 127 qubits) and a simulator, see Figure 11(b).
 An error violating $r_{3}^{\mathrm{p}}$ can be seen in the bottom left image that results from hardware imperfections.
-Second, images from a $10\times 10\times 5$ voxel grid, for which QWFC is performed on a simulator, see fig.
-11 (c).
-4 CONCLUSIONS
-In this work, QWFC was proposed as a method to realize WFC on a quantum computer for PCG.
-The idea is to construct a quantum circuit from which valid content instance (that fulfill the prescribed patterns) can be sampled via measurements.
-In other words, the quantum circuit acts as a special-purpose random generator for content of a desired form that makes use of the intrinsic randomness of quantum physics.
-This method is not without its challenges.
-The corresponding quantum circuit becomes larger from both the content and alphabet size and at the same time deeper from the correlations within the content instance distribution, which makes it difficult to evaluate most use cases on current hardware.
-For this reason, the quantum-classical hybrid method HWFC was proposed that decomposes the problem into smaller sub-problems which can then be executed on the quantum computer with a higher chance of success.
+Second, images from a $10\times 10\times 5$ voxel grid, for which QWFC is performed on a simulator, see Figure 11(c).
+CONCLUSION
+In this work, QWFC was proposed as a method to realize WFC on a gate-based quantum computer as an example of QPCG.
+The idea is to construct a quantum circuit from which valid content instances (that fulfill the prescribed patterns) can be sampled via measurements.
+In other words, the quantum circuit acts as a special-purpose QRNG for content of a desired form that makes use of the intrinsic randomness of quantum physics.
+This means that the actual quantum mechanical collapse of the wave function is used for the implementation, from which the classical method derived its name.
+The proposed method is not without its challenges.
+The generating quantum circuit becomes larger from both the content and alphabet size and at the same time deeper from correlations within the content instance distribution, which makes it difficult to evaluate practical use cases on current hardware.
+For this reason, HWFC was proposed as a hybrid method that partitions the generation into smaller tasks for QWFC.
 The experimental results have shown the limitations of QWFC and HWFC, but have also proven that at least simple examples can already be implemented on today’s quantum hardware.
-Furthermore, all experiments were carried out without quantum error correction [17], which can be used to possibly improve the results.
 The biggest bottleneck of the circuit complexity is the conditional preparation of probability distributions.
-Investigating possible improvements in this area with regard to specific PCG use cases or alternative preparation approaches (e.g., qGANs [18]) could therefore be a potential direction for future research.
-The conceptual advantage of QWFC is that a circuit, once prepared, represents the distribution of content instances.
-This means that new instances can be generated by simple measurements without any additional algorithmic effort.
-A shortcoming of QWFC is that only a fixed order of segments is considered instead of a dynamic (e.g., entropy-based) order that depends on the previously selected segments.
-The realization of such a generalization of the method could also be a point of origin for further studies.
-CWFC is a strong method with many areas of application.
-In this paper, only visual content creation use cases were presented, but QWFC can also be used for non-visual contents like text [9] in analogy to CWFC [19].
-In conclusion, the paper aims to provide an initial approach to QWFC that demonstrates its feasibility but leaves room for improvement and raises the question of useful application areas given the limits of NISQ devices.
-5 ACKNOWLEDGMENTS
+Investigating possible improvements in this area with regard to specific QPCG use cases could therefore be a potential direction for future research.
+The conceptual advantage of QWFC is that a circuit, once designed, represents the distribution of content instances such that new instances can be generated by simple measurements without any additional algorithmic effort.
+A shortcoming of QWFC is that only a fixed order of segments is considered instead of a dynamic (e.g., entropy-based) order that depends on the previously selected segments, which is a chance for future improvements.
+In conclusion, the paper aims to provide an initial approach to QPCG that demonstrates its feasibility but leaves room for improvement and raises the question of useful application areas given the limits of NISQ devices.
+APPENDIX
+This appendix section contains the technical details of the presented PCG methods.
+Technical Details of PIPCG
+Presume the generated partial content instance $C(k)\in\mathcal{C}(k)$ for $k\in[1,N]$, where $\mathcal{C}(k)$ denotes all possible partial content instances at the end of iteration $k$ with $\mathcal{C}(N)\equiv\mathcal{C}$.
+One has $|C(k)|=k$, $C(k)=C(k-1)\cup\{(s(k),v_{s(k)})\}$, $s(i)\neq s(j)\,\forall\,i,j\in[1,k]\land i\neq j$, and $|\mathcal{C}(k)|=\sum_{I\in J(k)}W^{|I|}$, where $J(k):=\{I\,|\,I\in[1,N]\land|I|=k\}$ with $|J(k)|={N\choose k}$.
+Let $C(0):=\{\}$ and $\mathcal{C}(0):=\{C(0)\}$.
+Furthermore, $S^{+}(C(l)):=\{s(1),\dots,s(l)\}$ with $|S^{+}(C(l))|=l$ and $S^{-}(C(l)):=[1,N]\setminus S^{+}(C(l))$ with $|S^{-}(C(l))|=N-l$ for all $C(l)\in\mathcal{C}(l)$ and for all $l\in[0,N]$.
+In the following, the three steps within each iteration $k$ of PIPCG are explained. In the first step, an identifier is selected by drawing a sample $s(k)$ from the random variable
+$\boldsymbol{s}(k)\sim p^{\mathrm{s}}_{k}(i=s(k)|C^{\prime}=C(k-1)).$ (7)
+The probability distribution $p_{k}^{\mathrm{s}}(i|C^{\prime})$ with support $i\in[1,N]$ and $C^{\prime}\in\mathcal{C}(k-1)$ is a user-defined parameter of the procedure that determines how new segment indices $k$ are selected based on the partial content instance $C^{\prime}=C(k-1)$ that has already been generated up to the previous iteration $k-1$ (making PIPCG non-Markovian).
+Two conditions must hold:
+  1. S1)
+    No identifier can be chosen twice, i.e., $p_{l}^{\mathrm{s}}(i=s|C^{\prime}=C(l-1))=0\,\forall\,s\in S^{+}(C(l-1))\,\forall\,C(l-1)\in\mathcal{C}(l-1)\,\forall\,l\in[1,N]$.
+  2. S2)
+    At least one valid identifier has to be available, i.e., $\forall\,l\in[1,N]\,\forall\,C(l-1)\in\mathcal{C}(l-1)\,\exists\,s\in S^{-}(C(l-1)):p_{l}^{\mathrm{s}}(i=s|C^{\prime}=C(l-1))>0$.
+In the second step, a value $v_{s(k)}$ is selected for the segment by drawing a sample $v_{s(k)}$ from the random variable
+$\boldsymbol{v}_{s(k)}\sim p^{\mathrm{v}}_{k}(\nu=v_{s(k)}|i=s(k),C^{\prime}=C(k-1)).$ (8)
+The probability distribution $p^{\mathrm{v}}_{k}(\nu|i,C^{\prime})$ with support $\nu\in A$, $i\in[1,N]$ and $C^{\prime}\in\mathcal{C}(k-1)$ is a user-defined parameter of the procedure that determines how new values are selected based on the selected identifier $i=s(k)$ of the current iteration and $C^{\prime}=C(k-1)$.
+Two conditions must hold:
+  1. V1)
+    Only values from the alphabet are available, i.e., $p^{\mathrm{v}}_{l}(\nu=v|i=s,C^{\prime}=C(l))=0\,\forall\,v\not\in A\,\forall\,s\in[1,N]\,\forall\,C(l)\in\mathcal{C}(l-1)\,\forall\,l\in[1,N]$.
+  2. V2)
+    At least one value from the alphabet has to be available, i.e., $\forall\,l\in[1,N]\,\forall\,C(l-1)\in\mathcal{C}(l-1)\forall s\in S^{-}(C(l-1))\exists v\in A:p_{l}^{\mathrm{v}}(\nu=v|i=s,C^{\prime}=C(l))>0$.
+In the third step, $C(k)=C(k-1)\cup\{(s(k),v_{s(k)})\}$. The process is repeated for $k\rightarrow k+1$ until $k=N$.
+The probability to generate $C(k)$ is given by $p(C(k)):=\sum_{\sigma\in S_{k}(C(k))}\prod_{l=1}^{k}q_{l}^{k}(\sigma,C(k))$, where
+$q_{l}^{k}(\sigma,C(k)):=p_{l}($ $i=\sigma_{l},\nu=v_{\sigma_{l}}|$
+$C^{\prime}\!=\!\{(\sigma_{1},v_{\sigma_{1}}),\!\dots,\!(\sigma_{l-1},v_{\sigma_{l-1}})\}).$ (9)
+with $p_{l}(i,\nu|C^{\prime}):=p_{k}^{\mathrm{s}}(i|C^{\prime})p_{k}^{\mathrm{v}}(\nu|i,C^{\prime})$, $S_{k}(C(k))$ as the set of all permutations of $S^{+}(C(k))$ with $|S_{k}(C(k))|=k!$, and the $l$th value of a permutation $\sigma\in S_{k}(C(k))$ denoted by $\sigma_{l}$ for $l\in[1,k]$.
+In Technical Details of PIPCG, let $C^{\prime}=\{\}$ for $l=1$.
+$v_{\sigma_{m}}$ represents the value of the segment with identifier $\sigma_{m}$ from the content instance $C(k)$ for $m\in[1,l]$.
+The support $\underline{\mathcal{C}}(k):=\{C(k)\,|\,C(k)\in\mathcal{C}(k)\land p(C(k))>0\}$ of $p(C(k))$ contains all $C(k)$ that can be generated for the given $k$.
+For $k=N$, Equation 2 emerges with $p(v_{1},\dots,v_{N}):=\sum_{\sigma\in S_{N}}\prod_{l=1}^{N}q_{l}^{N}(\sigma,C)$ and $\underline{\mathcal{C}}(N)=\underline{\mathcal{C}}$.
+Technical Details of CWFC
+For the value selection, a general rule-based selection mechanism is introduced first and then the pattern-based selection is presented as a special case.
+Consider a set of $m$ rules $R\in\mathcal{R}$, Equation 5, with the set of of all possible rulesets $\mathcal{R}$.
+Each rule $r_{i}:=r_{i}(v,w):=(v,w)$ for $i\in[1,m]$ consists of a value $v\in A$ and a weight function $w:=w(i,\alpha^{d},C^{\prime})\in\mathbb{R}_{\geq 0}$ with $w:[1,N]\times\mathcal{B}\times\overline{\mathcal{C}}\rightarrow\mathbb{R}_{\geq 0}$, where $\overline{\mathcal{C}}:=\cup_{k=0}^{N-1}\mathcal{C}(k)$ and :=$\mathcal{B}:=\{0,1\}^{N\times N}$.
+Given $R\in\mathcal{R}$, $p_{k}^{\mathrm{v}}(\nu|i,C^{\prime})=p_{k}^{\mathrm{v,R}}(\nu|i,C^{\prime}):=\frac{F_{k}(\nu,i,C^{\prime})}{\overline{F}_{k}(i,C^{\prime})}$, Equation 8, with $F_{k}(\nu,i,C^{\prime}):=\sum_{\begin{subarray}{c}(v,w)\in R\\ v=\nu\end{subarray}}w(i,\alpha^{d},C^{\prime})$ and $\overline{F}_{k}(i,C^{\prime}):=\sum_{\nu\in A}F_{k}(\nu,i,C^{\prime})$ for $\nu\in A$, $i\in[1,N]$, $C^{\prime}\in\mathcal{C}(k)$ for $k\in[1,N]$, and $\alpha^{d}\in\mathcal{B}$ for $d\in[1,D]$.
+Condition V1 is always satisfied, whereas in case of $\overline{F}_{k}(i,C^{\prime})\overset{!}{=}0$ for any $C^{\prime}\in\mathcal{C}(k)$ with $p_{k}^{\mathrm{s,H}}(i|C^{\prime})p(C^{\prime})>0$, condition V2 is violated.
+Pattern-based value selection is a special case of rule-based value selection, where $w(i,\alpha^{d},C^{\prime})\!:=\!w^{\mathrm{P}}(i,\alpha^{d},C^{\prime},P,u)\!:=\!u\tau(i,\alpha^{d},C^{\prime},P)$ of each rule $r_{i}:=r_{i}^{\mathrm{P}}(v,P,u):=(v,w^{\mathrm{P}}(i,\alpha^{d},C^{\prime},P,u))$ for $i\in[1,m]$ is defined by $u\in\mathbb{R}_{>0}$ and $P\in\mathcal{P}$, Equation 3 with $d_{i}\in[1,D]$ for $i\in[1,n]$, $d_{i}\neq d_{j}\,\forall\,i,j\in[1,n]\land i\neq j$, and $v_{i}\in A$ for $i\in[1,n]$.
+The set of all possible patterns reads $\mathcal{P}:=\{P\,|\,P\in([1,D]\times A)^{n}\,\forall\,n\in[0,D]\}$.
+A pattern $P$ is fulfilled, if $\tau(i,\alpha^{d},C^{\prime},P):=\prod_{(d,v)\in P}\prod_{(s,v^{\prime})\in C^{\prime}\land\alpha^{d}_{i,s}=1}\delta_{vv^{\prime}}\in[0,1]$ evaluates to one.
+The set of all possible pattern-based rulesets is denoted by $\mathcal{R}^{\mathrm{P}}\subset\mathcal{R}$.
+Summarized, the weight function $w(i,\alpha^{d},C^{\prime})$, yields $u$ if the pattern $P$ applies and zero otherwise.
+Only adjacent segments (i.e., $\alpha^{d}_{i,s}=1$) have an influence on $p_{k}^{\mathrm{v}}(\nu|i,C^{\prime})$.
+The rules $r_{i}$ can be rewritten in form of Equation 4.
+One has $p_{k}^{\mathrm{v,R}}(\nu|i,C^{\prime})=p_{k}^{\mathrm{v,P}}(\nu|i,C^{\prime}):=\sum_{\begin{subarray}{c}(v,u,P)\in R\\ v=\nu\end{subarray}}\gamma(u,i,C^{\prime})\tau(i,\alpha^{d},C^{\prime},P)$ with $\gamma(u,i,C^{\prime}):=\frac{u}{\overline{F}_{k}(i,C^{\prime})}\in[0,1]$.
+Optionally, a functional factor instead of a constant factor can be used, i.e., $u:=u(i,C):[1,N]\times\overline{\mathcal{C}}\rightarrow\mathbb{R}$.
+For the entropic segment identifier selection, one has
+$p_{k}^{\mathrm{s}}(i|C^{\prime})=p_{k}^{\mathrm{s,H}}(i|C^{\prime}):=\begin{cases}|I_{k}(C^{\prime})|^{-1}&\text{if }i\in I_{k}(C^{\prime})\\ 0&\text{otherwise}\end{cases}$ (10)
+in Equation 7 with $I_{k}(C^{\prime}):=\{i\,|\,i\in S^{-}(C^{\prime})\land H_{k}(i,C^{\prime})=h_{k}(C^{\prime})\}$ and $h_{k}(C^{\prime}):=\operatorname*{arg\,min}_{j\in S^{-}(C^{\prime})}H_{k}(j,C^{\prime})$, where $H_{k}(i,C^{\prime}):=-\sum_{v\in A}p_{k}^{\mathrm{v,P}}(\nu=v|i,C^{\prime})\ln p_{k}^{\mathrm{v,P}}(\nu=v|i,C^{\prime})$ denotes the Shannon information entropy for $C^{\prime}\in\mathcal{C}(k-1)$, $k\in[1,N]$.
+Technical Details of QWFC
+Presume a predefined order of segment indices $\sigma\in S_{N}$.
+Hence, $p_{k}^{\mathrm{s}}(i|C^{\prime})=p_{k}^{\mathrm{s,o}}(i|C^{\prime}):=\delta_{i\sigma_{i}}$ and $\mathcal{C}(k)=\mathcal{C}_{\sigma}(k):=\{C\,|\,C=\{(\sigma_{1},v_{\sigma_{1}}),\dots,(\sigma_{k},v_{\sigma_{k}})\}\,\forall\,v_{\sigma_{l}}\in A\,\forall\,l\in[1,k]\}$ for $k\in[1,N]$ and $\mathcal{C}_{\sigma}(0):=\{\{\}\}$.
+The value selection is pattern-based in analogy to CWFC, which together with the identifier selection leads to Equation 6.
+Given $\sigma$ and $R\in\mathcal{R}^{\mathrm{P}}$, the task is to prepare |$\ket{\Psi}$ from Equation 6 constructively with quantum gates.
+Let |$\ket{\Psi}:=U_{N}\cdots U_{1}\ket{0}$ as in Figure 4(a).
+Then,
+$U_{k}:=\left[U_{k}^{\mathrm{coin}}+U_{k}^{0}\right]\otimes\bigotimes_{l=k+1}^{N}\mathbbm{1}_{\sigma_{l}}\in\mathcal{H}$ (11)
+for $k\in[1,N]$, which contains $\mathbbm{1}_{i}\in\mathcal{H}_{i}$, the unit operator for $i\in[1,N]$ and two other operators.
+First, the conditional initialization operator
+$U_{k}^{\mathrm{coin}}=\sum_{C^{\prime}\in\mathcal{C}_{\sigma}(k)}\left[U_{k}^{\mathrm{ctrl}}(C^{\prime})\otimes U_{k}^{\mathrm{init}}(C^{\prime})\right],$ (12)
+where $U_{k}^{\mathrm{ctrl}}(C^{\prime}):=\bigotimes_{(s,v)\in L_{k}(C^{\prime})}\ket{v-1}\bra{v-1}_{s}\in\bigotimes_{l=1}^{k-1}\mathcal{H}_{\sigma_{l}}$ denotes the control operator, and $U_{k}^{\mathrm{init}}(C^{\prime}):=\ket{C^{\prime}}\bra{0}_{\sigma_{k}}\in\mathcal{H}_{\sigma_{k}}$ the initialization operator with |$\ket{C^{\prime}}_{\sigma_{k}}:=\sum_{\nu\in A}\sqrt{p_{k}^{\mathrm{v,P}}(\nu|\sigma_{k},C^{\prime})}\ket{\nu-1}_{\sigma_{k}}\in\mathcal{H}_{\sigma_{k}}$.
+The summation in $U_{k}^{\mathrm{ctrl}}(C^{\prime})$ runs over all adjacent segments $L_{k}(C^{\prime}):=\{(s,v)\,|\,(s,v)\in C^{\prime}\,\land\,\exists\,P\in R_{P}\exists\,d\in P_{d}:\alpha^{d}_{\sigma_{k}s}=1\}\subseteq C^{\prime}$, where $R_{P}:=\{P_{1},\dots,P_{m}\}$ stands for the set of patterns in $R\in\mathcal{R}^{\mathrm{P}}$ with $P_{i}\in\mathcal{P}$ for $i\in[1,m]$ and $P_{d}:=\{d_{1},\dots,d_{n}\}$ stands for the set of directions in the pattern $P\in\mathcal{P}$, Equation 3, with $d_{j}\in[1,D]$ for $j\in[1,n]$.
+Choosing $C^{\prime}$ instead of $L_{k}(C^{\prime})$ in the summation has no effect on $p_{k}^{\mathrm{v,P}}(\nu|i,C^{\prime})$ and thus |$\ket{C^{\prime}}_{\sigma_{k}}$.
+In iteration $k$, only the qubits in $\mathcal{Q}_{k-1}^{+}$ with $\mathcal{Q}_{k}^{+}:=\{q\,|\,q=q_{k}^{i}\in\mathcal{Q}_{\sigma_{k}}\,\forall\,i\in[1,q]\,\land\,\sigma_{k}\in L_{k}^{+}\}$ are affected by $U^{\mathrm{ctrl}}(C^{\prime})$ for some $C^{\prime}\in\mathcal{C}_{\sigma}(k)$, where $L_{k}^{+}:=\{s\,|\,\exists\,C^{\prime}\in\mathcal{C}_{\sigma}(k)\,:\,s\in S^{+}(L_{k}(C^{\prime}))\}$ denotes the affected segments for $k=[1,N]$.
+Here, $\mathcal{Q}_{k}^{+}\subseteq\{\mathcal{Q}_{1},\dots,\mathcal{Q}_{k}\}$ and $L_{k}^{+}\subseteq\{\sigma_{1},\dots,\sigma_{k}\}$, respectively.
+The second operator in Equation 11 is the neutral operator
+$U_{k}^{0}=\sum_{C^{\prime}\in\mathcal{C}_{\sigma}(k)}\left[U_{k}^{\mathrm{ctrl}}(C^{\prime})\otimes U_{k}^{\mathrm{init},0}+U_{k}^{\mathrm{ctrl},0}(C^{\prime})\right].$ (13)
+Initially, all qubits $\mathcal{Q}_{k}$ are in the ground state.
+Therefore, $U_{k}^{\mathrm{init},0}:=\sum_{v\in[2,W]}\ket{v-1}\bra{v-1}_{\sigma_{k}}\in\mathcal{H}_{\sigma_{k}}$ vanishes.
+Furthermore, $U_{k}^{\mathrm{ctrl},0}(C^{\prime}):=\left[\bigotimes_{l=1}^{k-1}\mathbbm{1}_{\sigma_{l}}-U_{k}^{\mathrm{ctrl}}(C^{\prime})\right]\otimes\mathbbm{1}_{\sigma_{k}}\in\bigotimes_{l=1}^{k}\mathcal{H}_{\sigma_{l}}$ does not change the state by construction.
+To realize $U_{k}$ with quantum gates for a given $k\in[1,N]$, only $U_{k}^{\mathrm{coin}}$ has to be considered (because all other parts of $U_{k}$ have no effect on the resulting state), which consists of a sequence of joint pairs $U_{k}^{\mathrm{ctrl}}(C^{\prime})\otimes U_{k}^{\mathrm{init}}(C^{\prime})$, one for each $C^{\prime}\in\mathcal{C}_{\sigma}(k)$.
+Each of these joint pairs represent a conditional loading of a probability distribution to realize |$\ket{C^{\prime}}_{\sigma_{k}}$, which is a well-known task that may, however, require an exponential number of gates [23].
+In the original WFC, a pattern conflict leads to a restart of the algorithm.
+This can also be realized in QWFC by adding an additional “conflict detection qubit” that stores this information, making S1, S2, V1, and V2 obsolete.
+This optional extension is not further discussed.
+Technical Details of HWFC
+Presume $H$ partitions such that partition $h\in[1,H]$ contains all of the segment identifiers from the set $S^{h}\subset[1,N]$ with $\bigcup_{h=1}^{H}S^{h}=[1,N]$ and $\bigcap_{h=1}^{H}S^{h}=\{\}$.
+The already sampled sub-contents $C^{1},\dots,C^{h-1}$ are used as an additional constraint, which has to be taken account in the quantum circuit by replacing $\mathcal{C}_{\sigma}(k)$ in Equations 12 and 13 with $\mathcal{C}^{h}_{\sigma}(k):=\{C\,|\,C=\bigcup_{i=1}^{h-1}C^{i}\cup C^{\prime}\,\forall\,C^{\prime}\in\mathcal{C}_{\sigma}(k)\}$.
+ACKNOWLEDGMENTS
 Parts of this research have been funded by the Ministry of Science and Health of the State of Rhineland-Palatinate (Germany) as part of the project AnQuC.